### v1

# Cosmological dynamics of interacting dark energy and dark matter in $f(Q)$ gravity

Abstract
In this work, we explore the behavior of interacting dark energy and dark matter within a model of $f(Q)$ gravity, employing a standard framework of dynamical system analysis. We consider the power-law $f(Q)$ model incorporating with two different forms of interacting dark energy and dark matter: $3\alpha H\rho_{m}$ and $\frac{\alpha}{3H}\rho_{m}\rho_{DE}$. The evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$ has been examined. Our results show that the universe was dominated by matter in the early stages and will be dominated by dark energy in later stages. Using the observational data, the fixed points are found to be stable and can be represented the de Sitter and quintessence acceleration solutions. We discover that the dynamical profiles of the universe in $f(Q)$ dark energy models are influenced by both the interaction term and the relevant model parameters.
Keywords: $f(Q)$ Gravity; Dark Energy; Dark Matter; Dynamical Analysis
June 4, 2024
I Introduction
Observational data has confirmed that the universe has entered a period of accelerated expansion. This evidence comes from various independent sources, including Supernovae Type Ia (SNe Ia) [1, 2, 3], temperature anisotropies in the Cosmic Microwave Background (CMB) observed by WMAP [4, 5], and Baryon Acoustic Oscillations [6, 7]. According to the framework of general relativity (GR), this accelerated expansion is driven by a new energy density component with negative pressure, referred to as dark energy (DE). Despite its crucial role, little is known about DE. Understanding the origin of DE, which causes the current cosmic acceleration, remains one of the major unresolved issues in modern cosmology. The standard equation of state $\omega=p/\rho$, where $\omega\geq 0$, cannot explain this phase of cosmic acceleration. Instead, an unknown component with negative pressure, characterized by $\omega<-\tfrac{1}{3}$, is necessary to describe the universe’s acceleration. However, the exact nature of DE is still a mystery. Numerous attempts have been made to address this significant problem, as discussed in Refs.[8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The most straightforward candidate for dark energy is the cosmological constant, $\Lambda$. However, as pointed out in Ref.[23], the cosmological constant is vastly smaller than what is predicted by modern elementary particle theories. To differentiate between the cosmological constant and other potential sources of dark energy that change over time, researchers use dynamical models to study the evolution of the equation of state of DE. Therefore, a mechanism must be identified to achieve a small value of $\Lambda$ that aligns with observations.
It is well established that the pure cosmological constant cannot explain the accelerated expansion of the early universe, as it cannot connect to the radiation-dominated era. An alternative and compelling candidate for both dark energy and inflation is a scalar field with a slowly varying potential. In the realm of dark energy, numerous scalar field models have been proposed, including quintessence [24, 25, 26, 27, 28, 29] and k-essence [30, 31, 32, 33] scenarios. These models predict various behaviors for the equation of state of dark energy. However, current observational data cannot distinguish these models from the $\Lambda$-cold-dark-matter ($\Lambda$CDM) model. Moreover, achieving viable scalar-field models within the framework of particle physics is challenging due to the extremely small scalar mass required for present-day cosmic acceleration [34, 35]. Another approach to explaining the universe’s accelerated expansion involves dynamical dark energy models based on modifying gravity over large distances. Examples of such models include $f(R)$ gravity, where the Ricci scalar is replaced by a more general function, $f(R)$ [36, 37, 38, 39], as well as scalar-tensor theories [40, 41, 42, 43, 44, 45], Galileon gravity [46], and Gauss-Bonnet gravity [47, 48]. For several decades, dynamical system analysis has been used in cosmology to qualitatively study these models, proving effective in identifying and classifying their asymptotic behaviors [49, 50]. Recently, this framework has been applied to investigate various dark energy models, including those based on $f(R)$ gravity [13, 51]. A series of papers [52, 54, 53, 55, 56, 57] has systematically analyzed $f(R)$ theories, identifying models that provide correct cosmological evolution from a vast number of possibilities. These analyses have led to the construction of cosmologically viable $f(R)$ gravity models, which feature appropriate trajectories in the dynamical system phase space and are compatible with local gravity constraints [20, 21, 55]. A notable advantage of these models is that they achieve cosmic acceleration without requiring a dark energy component. Furthermore, these models impose stringent constraints from local gravity tests and various observational data, making them more rigorous compared to modified matter models.
The authors of Refs.[58, 59, 60, 61, 62] initiated and expanded the phase space analysis of $f(R)$ gravity using the autonomous dynamical systems approach. While there has been considerable research on dynamical systems in $f(R)$ gravity involving non-interacting dark energy and dark matter [51], the study of interacting dark energy within the $f(R)$ gravity framework has received less attention, particularly in the context of dynamical system analysis [63, 64, 65, 66, 67, 68]. Investigating interacting dark energy and dark matter is valuable for understanding the behavior of accelerated scaling solutions where the energy densities of dark energy ($\Omega_{DE}$) and dark matter ($\Omega_{DM}$) are approximately equal, and the universe is accelerating. This research could also elucidate the coupling between dark matter structure formation and the time evolution of dark energy, potentially explaining why dark energy dominates over dark matter in the late universe. Additionally, the impact of interacting dark energy models on the global 21cm signal was explored in Ref.[70], and some interacting dark energy models and their potential to cause future singularities have been analyzed in Ref.[71].
The purpose of this work is to investigate the interaction between dark energy and dark matter within viable models of $f(Q)$ gravity. The $f(Q)$ gravity theory [72, 73] is one of the most promising approaches to explaining the accelerated expansion of the universe, involving dynamical dark energy models that modify gravity at large distances. This theory extends the alternative to General Relativity known as the symmetric teleparallel theory, which is based on the nonmetricity scalar $Q$, with both curvature and torsion being absent. Geometrically, the nonmetricity $Q$ describes the variation in the length of a vector during parallel transport. $f(Q)$ gravity presents intriguing applications [74, 75, 76, 77, 78, 79] and easily satisfies the constraints from gravitational wave observations [80]. At cosmological scales, $f(Q)$ can create a gravitational interaction that reveals new and interesting patterns in both the background evolution [72, 73, 81, 82, 83, 84] and the propagation of scalar and tensor perturbations [73, 81, 74, 85, 86, 87]. Additionally, certain forms of the $f(Q)$ function have been shown to mitigate the $\sigma_{8}$ tension [88], while others provide a better fit to cosmological data [89, 90, 91, 92, 93, 94, 95, 96, 97]. This theory is thus challenging $\Lambda$CDM from different perspectives [98]. Furthermore, a dynamical system analysis of both background and perturbation equations is conducted in $f(Q)$ gravity to independently verify the validity of result [87, 99, 100, 101, 102, 103, 104].
The structure of this work is organized as follows: In Section II, we present the fundamental cosmological equations of the general $f(Q)$ theory and derive the Friedmann equation corresponding to the FLRW metric. In Section III, we derive the autonomous dynamical system within the framework of $f(Q)$ gravity theory, focusing on the interaction between dark energy and dark matter. In Section IV, we apply the power-law $f(Q)$ model to enclose the dynamical system and conduct a further study of $f(Q)$ gravity. Finally, we discuss our findings in Section V.
II Formulation of $f(Q)$ gravity theory
In this section, we briefly discuss the formulation of $f(Q)$ gravity, a simple generalization of symmetric teleparallel gravity theory. This theory requires a different curvature and torsion-free connection, i.e., it is wholly dependent on nonmetricity. This formulation introduces new dynamics compared to General Relativity (GR) by allowing the function $f(Q)$ to dictate how nonmetricity influences the gravitational interaction. The non-metricity tensor $Q_{\sigma\mu\nu}$ is defined as
$Q_{\sigma\mu\nu}=\nabla_{\sigma}g_{\mu\nu},$ (1)
which geometrically describes the variation of the length of a vector in the parallel transport.
With the aid of the nonmetricity tensor, we can define the superpotential tensor $P_{\,\,\mu\nu}^{\sigma}$ as
$4P_{\,\,\mu\nu}^{\sigma}=-Q^{\sigma}_{\,\,\,\,\mu\nu}+2Q^{\,\,\,\,\,\,\sigma}_{(\mu\,\,\,\,\nu)}-Q^{\sigma}g_{\mu\nu}-\tilde{Q}^{\sigma}g_{\mu\nu}-\delta^{\sigma}_{(\mu}\,Q\,_{\nu)},$ (2)
where $Q_{\sigma}=Q_{\sigma\,\,\,\,\mu}^{\,\,\,\,\mu}$ and $\tilde{Q}_{\sigma}=Q^{\mu}_{\,\,\,\,\sigma\mu}\,$.
From the above quantities, we can obtain a non-metricity scalar as
$Q=-Q_{\sigma\mu\nu}P^{\sigma\mu\nu}.$ (3)
The action for $f(Q)$ gravity [72]
$S=\int\left[\frac{1}{2\kappa^{2}}f(Q)+\mathcal{L}_{m}\right]\sqrt{-g}\,d^{4}x,$ (4)
where $f(Q)$ represents any function of the scalar $Q$, $g$ denotes the determinant of $g_{\mu\nu}$, and $\mathcal{L}_{m}$ stands for the matter Lagrangian density.
The equations of motion in $f(Q)$ gravity are derived by varying the action with respect to the metric and the connection. This leads to a set of modified field equations that incorporate the effects of nonmetricity, providing a richer structure for modeling gravitational phenomena, and it is written as
$\frac{2}{\sqrt{-g}}\nabla_{\sigma}\left(f_{Q}\sqrt{-g}\,P^{\sigma}_{\,\,\mu\nu}\right)+\frac{1}{2}f(Q)\,g_{\mu\nu}\\ +f_{Q}\left(P_{\mu\sigma\lambda}Q_{\nu}^{\,\,\,\sigma\lambda}-2Q_{\sigma\lambda\mu}P^{\sigma\lambda}_{\,\,\,\,\,\,\nu}\right)=-T_{\mu\nu},$ (5)
where $f_{Q}=\frac{df}{dQ}$. The energy-momentum tensor for matter is now defined as $T_{\mu\nu}\equiv-\frac{2}{\sqrt{-g}}\frac{\delta(\sqrt{-g})\mathcal{L}_{m}}{\delta g^{\mu\nu}}$.
To apply $f(Q)$ gravity in a cosmological context, we consider the spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) spacetime, characterized by the metric
$ds^{2}=-dt^{2}+a^{2}(t)\,\delta_{ij}\,dx^{i}\,dx^{j},\,\,\,\,\,\,(i,j=1,2,3),$ (6)
where $a(t)$ is the cosmological scale factor. For this metric, the corresponding nonmetricity scalar is given by $Q=6H^{2}$, with $H=\frac{\dot{a}}{a}$ representing the Hubble parameter and the dot denotes a derivative with respect to the coordinate time $t$. Applying the FLRW metric into the general field equation (5), the Friedman equations of $f(Q)$ cosmology read as
$6H^{2}f_{Q}-\frac{f}{2}$ $=$ $\rho$ (7)
$\left(12H^{2}\,f_{QQ}+f_{Q}\right)\dot{H}$ $=$ $-\frac{1}{2}(p+\rho),$ (8)
where $f_{Q}=\frac{df}{dQ}$, and $f_{QQ}=\frac{d^{2}f}{dQ^{2}}$.
To study the interaction between dark energy and dark matter, we have to modify the conservation equations of the dark energy and the matter given above by adding some coupling term $\mathcal{U}$ as [68]
$\dot{\rho}_{DE}+3H(p_{DE}+\rho_{DE})=\mathcal{U}\,,$
$\dot{\rho}_{m}+3H\,\rho_{m}=-\mathcal{U}\,,$
$\dot{\rho}_{r}+4H\,\rho_{r}=0\,.$
In addition, the coupling term $\mathcal{U}$ between dark energy and dark matter can be interpreted as the exchange rate of energy density between these two components. When $\mathcal{U}>0$, energy is transferred from dark matter to dark energy, whereas when $\mathcal{U}<0$, energy is transferred from dark energy to dark matter.
III Autonomous dynamical system of interacting dark energy and dark matter
In this section, we derive the autonomous dynamical system within the framework of $f(Q)$ gravity theory, focusing on the interaction between dark energy and dark matter. Our approach involves incorporating two interacting terms. The first term relies solely on the energy density of the matter sector, the Hubble parameter, and a coupling constant $\alpha$, structured multiplicatively as $\mathcal{U}=3\alpha\,H\,\rho_{m}$. The second term encompasses the multiplication of energy densities from both sectors, capturing the immediate effects of both dark matter and dark energy on the interaction term, with a dimensionless coupling constant $\alpha$, denoted as $\mathcal{U}=\frac{\alpha}{3H}\rho_{m}\,\rho_{de}$.
According to the Friedmann equation (7), we can define the dimensionless variable as
$x=\frac{f}{12H^{2}\,F},\,\,\,\,\,y=\frac{\rho_{r}}{6H^{2}\,F}.$ (9)
Invoking the above dimensionless variables, it is straightforward to derive the autonomous equations which play a key role for studying the dynamical system for the interacting dark energy and dark matter.
III.1 Case I : $\mathcal{U}=3\alpha\,H\,\rho_{m}$
Using Eq.(7), the autonomous dynamical system is given by
$\frac{dx}{dN}=(1-x)\frac{\dot{H}}{H^{2}}+3x-3x^{2}+xy+3\alpha\,x(1-x-y),$ (10)
$\frac{dy}{dN}=-\frac{\dot{H}}{H^{2}}\,y-y+y^{2}-3xy+3\alpha\,y(1-x-y).$ (11)
where the variable $N=ln\,a$ is e-folding number and leads to $\frac{d}{dN}=\frac{1}{H}\frac{d}{dt}$.
III.2 Case II : $\mathcal{U}=\frac{\alpha}{3H}\rho_{m}\,\rho_{de}$
For this particular case, the autonomous dynamical system is given by
$\frac{dx}{dN}=(1-x)\frac{\dot{H}}{H^{2}}+3x-3x^{2}+xy+\alpha\,x^{2}(1-x-y),$ (12)
$\frac{dy}{dN}=-\frac{\dot{H}}{H^{2}}\,y-y+y^{2}-3xy+\alpha\,xy(1-x-y).$ (13)
where the variable $N=ln\,a$ is e-folding number and leads to $\frac{d}{dN}=\frac{1}{H}\frac{d}{dt}$.
Furthermore, we obtain the generalized equation for $\frac{\dot{H}}{H^{2}}$ as
$\frac{\dot{H}}{H^{2}}=\frac{-3f_{Q}(1-x+\frac{y}{3})}{2Qf_{QQ}+f_{Q}}.$ (14)
Consequently, the density parameters for individual matter species can be linked to the dimensionless variables outlined in Eq. (9). These variables serve as a set of constraint equations:
$\Omega_{m}=1-x-y,\,\,\,\,\,\,\,\Omega_{r}=y,\,\,\,\,\,\,\,\Omega_{DE}=x.$ (15)
Finally, we define the equation of state and deceleration parameters corresponding to the dimensionless variables to check the appropriate acceleration expansion of the universe, that is
$\omega=-1+\frac{2f_{Q}(1-x+\frac{y}{3})}{2Qf_{QQ}+f_{Q}}$ (16)
and
$q=-1+\frac{3f_{Q}(1-x+\frac{y}{3})}{2Qf_{QQ}+f_{Q}}.$ (17)
Both $\omega$ and $q$ represent different aspects of cosmic evolution, each uniquely influencing the large-scale structure of the universe. Grasping the significance of these parameters is essential for studying the effects of dark energy across various stages of cosmic evolution. Moreover, these parameters aid in comparing and differentiating between various dark energy models, each employing distinct mechanisms to drive cosmic acceleration.
IV Model:- $f(Q)=6\gamma\,H_{0}^{2}\left(\frac{Q}{Q_{0}}\right)^{n}$
In the previous section, we derived autonomous dynamical systems, and our primary objective is to investigate and analyze them. To accomplish this goal, we will focus on the designated power-law $f(Q)$ model, which has the form $f(Q)=6\gamma\,H_{0}^{2}\left(\frac{Q}{Q_{0}}\right)^{n}$, where $\gamma$ and $n$ are free model parameters.
Corresponding to our power-law $f(Q)$ model, Eq. (14) can be simplified to obtain
$\frac{\dot{H}}{H^{2}}=-\frac{3(1-x+\frac{1}{3}y)}{2n-1}.$ (18)
Now, we will find the critical points and corresponding eigenvalues for both the dynamical systems. Studying the critical points and assessing their stability is crucial for a comprehensive understanding of critical aspects in cosmic evolution driven by dark energy and dark matter interactions, as explored within the scope of this study. The properties of the dynamical system depend on the values of the constants $\alpha$ and $n$.
IV.1 For case I
The fixed points $(x,y)$ of the general dynamical system are the following:
$P_{1}:(1,0),\,\,\,\,P_{2}:\left(\frac{1}{2n},\frac{2n-1}{2n}\right),$ (19)
$P_{3}:\left(\frac{1}{2n+\alpha(2n-1)},0\right).$ (20)
Furthermore, the corresponding eigenvalues for each fixed point are given by
$P_{1}:\{-4,-3(\alpha+1)\},\,\,\,\,P_{2}:\{4,1-3\alpha\},$ (21)
$P_{3}:\{3(\alpha+1),3\alpha-1\}.$ (22)
Let us now focus on the dynamics of each critical point and its features in the further subsections.
IV.1.1 $P_{1}$: the de Sitter fixed point
We start the discussion with the first fixed point. Here we have
$P_{1}:(x,y)=(1,0).$
This critical point is independent of both the model and coupling parameters. Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters are
$\Omega_{m}=0,\,\,\,\,\Omega_{DE}=1,\,\,\,\,\omega=-1.$ (23)
At this point, both dark matter and radiation are absent, and the universe is dominated by dark energy. The equation of state (EoS) at this fixed point indicates that the universe is experiencing an accelerated expansion. To study the stability of this fixed point, we can analyze the eigenvalues of the Jacobian matrix, which describe the behavior of the fixed point. The eigenvalues for the fixed point $P_{1}$ read
$\{-4,\,\,\,-3(\alpha+1)\}.$ (24)
The stability behavior of this fixed point (which depends on $\alpha$) is obtained as:
  • Stable node for $\alpha>-1$,
  • Saddle for $\alpha<-1$,
  • Non-hyperbolic for $\alpha=-1$.
IV.1.2 $P_{2}$: Non-metricity dominated fixed point
Next, for the second fixed point, we have
$P_{2}:(x,y)=\left(\frac{1}{2n},\frac{2n-1}{2n}\right).$
The characteristics of this critical point solely depend on the model parameter $n$. Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters are obtained as
$\Omega_{m}=0,\,\,\,\,\Omega_{DE}=\frac{1}{2n},\,\,\,\,\omega=-1+\frac{4}{3n}.$ (25)
The conditions for an accelerating universe are $n<0$ or $n>2$. When $n<0$, the universe exhibits phantom-like behavior, i.e., $\omega<-1$. When $n>2$, the universe exhibits quintessence-like behavior, i.e., $-1<\omega<-1/3$. In this fixed-point solution, a matter-dominated era occurs for $n=4/3$ and a radiation-dominated era for $n=1$. The eigenvalues for the fixed point $P_{2}$ are as follows:
{4,   $\{4,\,\,\,1-3\alpha\}.$ (26)
The stability behavior for this particular fixed point can be obtained for different $\alpha$, given by
  • Saddle node for $\alpha>1/3$,
  • Unstable for $\alpha<1/3$,
  • Non-hyperbolic for $\alpha=1/3$.
IV.1.3 $P_{3}$: scaling solution fixed point
It is worth noting that our result modifies the scaling solution of $f(Q)$ gravity with the interacting dark energy parameter $\alpha$. The fixed point $P_{3}$ reads
$P_{3}:(x,y)=\left(\frac{1}{2n+\alpha(2n-1)},0\right).$
The properties of this critical point are influenced by both the model parameter $n$ and the coupling parameter $\alpha$. Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters read as
$\Omega_{m}=\frac{1}{\alpha-2(\alpha+1)n}+1,\,\,\,\,\Omega_{DE}=\frac{1}{2n+\alpha(2n-1)},$
and
$\omega=\frac{3\alpha-2(\alpha+1)n+2}{2(\alpha+1)n-\alpha}.$
The conditions for an accelerating universe are shown in Fig. 1.
Figure 1: In this figure, we show the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe. The light blue region represents the quintessence-like behavior of an accelerating universe (i.e., $-1<\omega<-1/3$), while the golden region represents the phantom-like behavior of an accelerating universe (i.e., $\omega<-1$).
In this fixed point solution, we obtained a matter-dominated era for $n=\frac{3\alpha+2}{2(\alpha+1)}$, $1+\alpha\neq 0$, and radiation-dominated era for $n=\frac{5\alpha+3}{4(\alpha+1)}$, $1+\alpha\neq 0$.
The eigenvalues for the fixed point $P_{3}$ read as
[…]
We determine the stability of this fixed point for various ranges of $\alpha$. The ranges are as follows:
  • Stable node for $\alpha<-1$
  • Saddle for $-1<\alpha<1/3$
  • Unstable node for $\alpha>1/3$.
In Figs. 2 and 3, we illustrate the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$. This is done using the numerical solution of the dynamical system with the initial conditions $x(0)=0.7$ and $y(0)=0.00005$. The motivation behind the choice of initial conditions is as follows: $x$ represents the dark energy density, which currently has a value of $0.7$. Similarly, $y$ represents the radiation density, which currently has a value of $0.00005$.
Figure 2: The plot on the left displays density parameters for matter ($\Omega_{m}$), dark energy ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$. The right panel shows the deceleration parameter $q$ and the equation of state parameter $\omega$ as functions of redshift $z$. These two panels illustrate the trajectory for the positive coupling parameter, which can transfer energy from dark matter to dark energy.
Figure 3: The left panel displays density parameters for matter ($\Omega_{m}$), dark energy ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$. The right panel shows the deceleration parameter $q$ and the equation of state parameter $\omega$ as functions of redshift $z$. These two panels illustrate the trajectory for the negative coupling parameter, which can transfer energy from dark energy to dark matter.
For $n=3/2$ and $\alpha=1/2$ in Fig. 2, $\alpha>0$ indicates that the coupling term $\mathcal{U}>0$, signifying energy transfer from dark matter to dark energy. This figure shows that the universe was dominated by matter in the early stages and will be dominated by dark energy in later stages. Currently, the universe is dominated by dark energy, with parameter values $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, and $\Omega_{DE}=0.7$, $q_{0}=-0.55$, and $\omega_{0}=-0.70$. For these values, fixed point $P_{1}$ is stable and represents the de Sitter acceleration solution, while fixed point $P_{2}$ is a saddle-node, and $P_{3}$ is an unstable node that cannot demonstrate universal acceleration.
In Fig. 3, with $n=3/2$ and $\alpha=-2$, $\alpha<0$ indicates that the coupling term $\mathcal{U}<0$, that means energy transfers from dark energy to dark matter. This figure shows that the universe was dominated by dark energy in the early stages and will be dominated by dark matter at later stages. Currently, the universe remains dominated by dark energy, with parameter values $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, and $\Omega_{DE}=0.7$, $q_{0}=-0.563$, and $\omega_{0}=-0.71$. Here, fixed point $P_{3}$ is stable and exhibits acceleration for the early and present universe but fails to show acceleration for late times, while fixed point $P_{1}$ is a saddle-node, and $P_{2}$ is an unstable node. Fig. 4 shows the phase space trajectories of the fixed points $P_{1}$, $P_{2}$, and $P_{3}$.
Figure 4: Phase plot of $x$ vs $y$. The upper phase plot shows the stable behavior of $P_{1}$ for $\alpha>0$, and the lower phase plot shows the stable behavior of $P_{3}$ for $\alpha<-1$.
IV.2 For case II
Next, the second case in $f(Q)$ gravity is also investigated in detail for the interacting dark energy system. The models have been systematically constructed with possible ranges of their model parameters derived from the dynamical system analysis within the interacting dark energy framework. It is important to examine how the interaction affects the cosmological viability of the $f(Q)$ gravity models.
The fixed points $(x,y)$ of the dynamical system are the following:
$R_{1}:(1,0),\,\,\,\,R_{2}:\left(\frac{1}{2n},\frac{2n-1}{2n}\right),R_{3}:\left(\frac{1}{\alpha},\frac{-3\alpha+8n-1}{\alpha}\right),$
$R_{4}:\left(\frac{-\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{2\alpha n-\alpha},0\right),$ (28)
$R_{5}:\left(\frac{\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{2\alpha n-\alpha},0\right).$
The corresponding eigenvalues are
$R_{1}:\{-4,-3-\alpha\},\,\,\,\,R_{2}:\{4,1-\frac{\alpha}{2n}\},$
$R_{3}:\{\mathcal{A}_{1}-\frac{4n}{\alpha}+4,-\mathcal{A}_{1}-\frac{4n}{\alpha}+4\},$
$R_{4}:\left\{\frac{5n-1+\mathcal{A}_{2}}{1-2n},\frac{-2\mathcal{A}_{2}^{2}-2(3n-\alpha+2n\alpha)\mathcal{A}_{2}}{(2n-1)^{2}\alpha}\right\},$
$R_{5}:\left\{\frac{-2\mathcal{A}_{2}^{2}+2(3n-\alpha+2n\alpha)\mathcal{A}_{2}}{(2n-1)^{2}\alpha},\frac{1-5n+\mathcal{A}_{2}}{2n-1}\right\}.$
where $\mathcal{A}_{1}=\frac{2\sqrt{(2n-1)\left(3(\alpha-1)\alpha^{2}+8n^{3}-4n^{2}-6(\alpha-1)\alpha n\right)}}{\alpha-2\alpha n}$ and $\mathcal{A}_{2}=\sqrt{9n^{2}-3\alpha+6n\alpha}$.
IV.2.1 $R_{1}$: the de Sitter fixed point
Corresponding to fixed point $R_{1}:(x,y)=(1,0)$, the matter density, dark energy density, and effective EoS parameters are
$\Omega_{m}=0,\,\,\,\,\Omega_{DE}=1,\,\,\,\,\omega=-1.$ (29)
At this point, both dark matter and radiation are absent, and the universe is dominated by dark energy. The EoS at this fixed point indicates that the universe is experiencing an accelerated expansion. To study the stability of this fixed point, we can analyze the eigenvalues of the Jacobian matrix, which describe the behavior of the fixed point. The eigenvalues for the fixed point $R_{1}$ read
$\{{-4,\,\,\,-\alpha-3\}}.$ (30)
The stability behavior of this fixed point is as follows:
  • Stable node for $\alpha>-3$,
  • Saddle for $\alpha<-3$,
  • Non-hyperbolic for $\alpha=-3$.
IV.2.2 $R_{2}$: Non-metricity dominated fixed point
The fixed point $R_{2}$ can be used to explain the late-time accelerating expansion of the universe in $f(Q)$ gravity. The point is given by
$R_{2}:(x,y)=\left(\frac{1}{2n},\frac{2n-1}{2n}\right)$
The characteristics of this critical point solely depend on the model parameter $n$. Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters are obtained as
$\Omega_{m}=0,\,\,\,\,\Omega_{DE}=\frac{1}{2n},\,\,\,\,\omega=-1+\frac{4}{3n}.$ (31)
The conditions for an accelerating universe are $n<0$ or $n>2$. When $n<0$, the universe exhibits phantom-like behavior, i.e., $\omega<-1$. When $n>2$, the universe exhibits quintessence-like behavior, i.e., $-1<\omega<-1/3$. In this fixed-point solution, we have obtained a matter-dominated era for $n=4/3$ and a radiation-dominated era for $n=1$.
The corresponding eigenvalues for the fixed point $R_{2}$ are as follows:
{4,  $\{{4,\,\,1-\frac{\alpha}{2n}\}}.$ (32)
We determine the stability of this fixed point for various ranges of $\alpha$. The ranges are as follows:
  • Saddle node for $\alpha>2n$,
  • Unstable for $\alpha<2n$,
  • Non-hyperbolic for $\alpha=2n$.
IV.2.3 $R_{3}$: Scaling solution point
The fixed point $R_{3}$ represents a scaling solution for the universe. This scaling solution makes the ratio $\Omega_{m}/\Omega_{DE}$ constant. The fixed point in this case is given by
$R_{3}:(x,y)=\left(\frac{1}{\alpha},\frac{-3\alpha+8n-1}{\alpha}\right).$
Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters read as
$\Omega_{m}=4-\frac{8n}{\alpha},\,\,\,\,\Omega_{DE}=\frac{1}{\alpha},\,\,\,\,\omega=\frac{8}{3\alpha}-1.$ (33)
The conditions for an accelerating universe are $\alpha<0$ or $\alpha>4$. When $\alpha<0$, the universe exhibits phantom-like behavior, i.e., $\omega<-1$. When $\alpha>4$, the universe exhibits quintessence-like behavior, i.e., $-1<\omega<-1/3$. In this fixed-point solution, we have obtained a matter-dominated era for $\alpha=8/3$ and a radiation-dominated era for $\alpha=2$. The stability conditions are shown in Fig. 5.
Figure 5: The region plot illustrates the stable, unstable, and saddle behaviors associated with the fixed point $R_{3}$. In this plot, the stable region is very small $\left(\frac{1}{8}<n\leq 0.4393\,\,\text{and}\,\,0<\alpha<\frac{1}{3}(8n-1)\right)$.
IV.2.4 $R_{4}$: Scaling solution point
The fixed point $R_{4}$ represents the scaling solution point of the universe. In addition, it is worth noting that our result modifies the scaling solution with the interacting dark energy parameter $\alpha$. The point $R_{4}$ is
$R_{4}:(x,y)=\left(\frac{-\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{2\alpha n-\alpha},0\right).$
Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters are obtained as
$\Omega_{m}=\frac{-\alpha+\sqrt{-3\alpha+9n^{2}+6\alpha n}+(2\alpha+3)n}{\alpha(2n-1)}\,,$
$\Omega_{DE}=\frac{-\sqrt{3}\sqrt{-\alpha+3n^{2}+2\alpha n}-3n}{2\alpha n-\alpha}\,,\,\,\text{and}$
$w=\frac{-3\alpha-4\alpha n^{2}+2\sqrt{-3\alpha+9n^{2}+6\alpha n}+(8\alpha+6)n}{\alpha(1-2n)^{2}}\,.$
One can note that all three parameters depend on the model parameter $n$ as well as the interacting parameter $\alpha$. Check appendix A for calculations. The conditions for an accelerating universe and the stability are shown in Figs. 6 and 7.
Figure 6: This plot shows the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe. The light blue region represents the quintessence-like behavior of an accelerating universe (i.e., $-1<\omega<-1/3$), while the golden region represents the phantom-like behavior of an accelerating universe (i.e., $\omega<-1$).
Figure 7: The region plot illustrates the stable, unstable, and saddle behaviors associated with the fixed point $R_{4}$.
IV.2.5 $R_{5}$: Additional scaling solution point
Corresponding to the fixed point $R_{5}$, that is
$R_{5}:(x,y)=\left(\frac{\sqrt{3}\sqrt{-\alpha+3n^{2}+2\alpha n}-3n}{2\alpha n-\alpha},0\right)$
the matter density, dark energy density, and effective EoS parameters are given by
$\Omega_{m}=\frac{3n-\sqrt{-3\alpha+9n^{2}+6\alpha n}}{\alpha(2n-1)}+1\,,$
$\Omega_{DE}=\frac{\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{\alpha(2n-1)}\,,\,\,$
$w=\frac{-3\alpha-4\alpha n^{2}-2\sqrt{-3\alpha+9n^{2}+6\alpha n}+(8\alpha+6)n}{\alpha(1-2n)^{2}}\,.$
The conditions for an accelerating universe and the stability are depicted in Figs. 8 and 9.
Figure 8: The plot show the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe. The light blue region represents the quintessence-like behavior of an accelerating universe (i.e., $-1<\omega<-1/3$), while the golden region represents the phantom-like behavior of an accelerating universe (i.e., $\omega<-1$).
Figure 9: The region plot illustrates the stable, unstable, and saddle behaviors associated with the fixed point $R_{5}$.
In Figures 10 and 11, we illustrate the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$. For $n=3/2$ and $\alpha=4$ in Figure 10, $\alpha>0$ indicates that the coupling term $\mathcal{U}>0$, signifying energy transfer from dark matter to dark energy. This figure shows that the universe was dominated by matter in the early stages and will be dominated by dark energy in later stages. Currently, the universe is dominated by dark energy, with parameter values $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, and $\Omega_{DE}=0.7$, $q_{0}=-0.548$, and $\omega_{0}=-0.696$. For these values, the fixed points $R_{1}$ and $R_{4}$ are stable and represent the de Sitter and quintessence acceleration solutions, respectively.
In Figure 11, with $n=3/2$ and $\alpha=-4$, $\alpha<0$ indicates that the coupling term $\mathcal{U}<0$, meaning energy transfers from dark energy to dark matter. This figure shows that the universe was dominated by dark energy in the early stages and will be dominated by dark matter in later stages. Currently, the universe remains dominated by dark energy, with parameter values $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, and $\Omega_{DE}=0.7$, $q_{0}=-0.55$, and $\omega_{0}=-0.70$. For these values, the fixed points $R_{4}$ and $R_{5}$ have imaginary values, and the remaining point cannot exhibit stable behavior.
Figure 10: The upper panel displays density parameters for matter ($\Omega_{m}$), dark energy ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$. The lower panel shows the deceleration parameter $q$ and the equation of state parameter $\omega$ as functions of redshift $z$. These two panels illustrate the trajectory for the positive coupling parameter, which can transfer energy from dark matter to dark energy.
Figure 11: The left panel displays density parameters for matter ($\Omega_{m}$), dark energy ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$. The right panel shows the deceleration parameter $q$ and the equation of state parameter $\omega$ as functions of redshift $z$. These two panels illustrate the trajectory for the negative coupling parameter, which can transfer energy from dark energy to dark matter.
Figure 12: Phase plot of $x$ vs $y$. The phase plot above shows the stable behavior of $R_{1}$ and $R_{4}$ for $\alpha>0$.
V Conclusion
In this work, we have explored the behavior of interacting dark energy and dark matter within $f(Q)$ gravity, employing a standard framework of dynamical system analysis. We have considered the power-law $f(Q)=6\gamma\,H_{0}^{2}\left(\tfrac{Q}{Q_{0}}\right)^{n}$ model incorporating with two different forms of interacting dark energy and dark matter: $3\alpha H\rho_{m}$ and $\frac{\alpha}{3H}\rho_{m}\rho_{DE}$. The parameter $\alpha$ in the interacting terms plays a crucial role in determining the viable conditions and estimating the transition from the matter-dominated era (saddle point) to the dark energy-dominated era (stable node) in viable gravity models at late times. As a result, we have discovered fixed points that can be represented as the late-time accelerating universe in $f(Q)$ gravity. For the form of ${\cal U}=3\alpha H\rho_{m}$, we have illustrated the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$. For $n=3/2$ and $\alpha=1/2$ and $\alpha>0$, we found that the coupling term $\mathcal{U}>0$, signifying energy transfer from dark matter to dark energy implying that the universe was dominated by matter in the early stages and would be dominated by dark energy in later stages. With the current data, fixed point $P_{1}$ is stable and represents the de Sitter acceleration solution, while fixed point $P_{2}$ is a saddle node, and $P_{3}$ is an unstable node that cannot demonstrate universal acceleration. Moreover, we have considered another situation of which $n=3/2$ and $\alpha=-2$, $\alpha<0$. We discovered for the coupling term $\mathcal{U}<0$ that energy transfers from dark energy to dark matter. In this case, the universe was dominated by dark energy in the early stages and will be dominated by dark matter in later stages. With the current data, fixed point $P_{3}$ is stable and exhibits acceleration for the early and present universe but fails to show acceleration for late times, while fixed point $P_{1}$ is a saddle node, and $P_{2}$ is an unstable node.
For the form of ${\cal U}=\frac{\alpha}{3H}\rho_{m}\rho_{DE}$, the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$ have been examined. We have considered $n=3/2$ and $\alpha=4$ ($\alpha>0$) and found that the coupling term $\mathcal{U}>0$, signifying energy transfer from dark matter to dark energy. Our results show that the universe was dominated by matter in the early stages and would be dominated by dark energy in later stages. Using the observational data, the fixed points $R_{1}$ and $R_{4}$ were stable and would represent the de Sitter and quintessence acceleration solutions. Additionally, for $n=3/2$ and $\alpha=-4$ ($\alpha<0$), we found that the coupling term $\mathcal{U}<0$, meaning energy transfers from dark energy to dark matter implying that the universe was dominated by dark energy in the early stages and would be dominated by dark matter in later stages. Employing the current data, the fixed points $R_{4}$ and $R_{5}$ have imaginary values, and the remaining point cannot exhibit stable behavior.
Although many viable $f(Q)$ models explain the dark energy problem in cosmology, our qualitative results from this work can serve as guidelines for more detailed studies. They can also be used as complementary constraints on viable $f(Q)$ models, alongside other cosmological constraints on $f(Q)$ theories. Our framework, based on the cosmological dynamics of interacting dark energy and dark matter in $f(Q)$ gravity, constitutes the natural template for beyond the standard gravity model, for example, Teleparallel gravity and Gauss-Bonnet gravity, or even more generalizations.
Data Availability Statement
There are no new data associated with this article.
Acknowledgments
GNG acknowledges University Grants Commission (UGC), New Delhi, India for awarding a Senior Research Fellowship (UGC-Ref. No.: 201610122060). PKS acknowledges Science and Engineering Research Board, Department of Science and Technology, Government of India for financial support to carry out Research project No.: CRG/2022/001847 and IUCAA, Pune, India for providing support through the visiting Associateship program.
Appendix A Appendix
The conditions of stability for the fixed point $R_{3}$ are as follows:
  • For Stable Node:
    $\frac{1}{8}<n\leq 0.4393\,\,\text{and}\,\,0<\alpha<\frac{1}{3}(8n-1).$ (34)
  • For Unstable Node:
    $n=0\,\,\text{and}\,\,\left(-\frac{1}{3}<\alpha<0\,\,\text{or}\,\,0<\alpha\leq 1\right),$ (35)
    $n<0\,\,\text{and}\,\,\left(\frac{1}{3}(8n-1)<\alpha<2n\right),$ (36)
    $0<n\leq\frac{1}{8}\,\,\text{and}\,\,\left(\frac{1}{3}(8n-1)<\alpha<0\right),$ (37)
    $n>\frac{1}{2}\,\,\text{and}\,\,2n<\alpha<\frac{1}{3}(8n-1).$ (38)
  • While the saddle point is otherwise.
The conditions of stability for the fixed point $R_{4}$ are as follows:
  • For Stable node:
    $0<n\leq\frac{1}{8}\,\,\text{and}\,\,0<\alpha<-\frac{3n^{2}}{2n-1},$ (39)
    $\frac{1}{8}<n<\frac{1}{5}\,\,\text{and}\,\,\frac{1}{3}(8n-1)<\alpha<-\frac{3n^{2}}{2n-1},$ (40)
    $\frac{1}{2}<n\leq 1\,\,\text{and}\,\,\alpha>0,$ (41)
    $n>1\,\,\text{and}\,\,\left(-\frac{3n^{2}}{2n-1}<\alpha<-3\,\,\text{or}\,\,\alpha>0\right).$ (42)
  • For Unstable node:
    $n\leq\frac{1}{8}\,\,\text{and}\,\,\alpha<\frac{1}{3}(8n-1),$ (43)
    $\frac{1}{8}<n<\frac{1}{2}\,\,\text{and}\,\,\alpha<0.$ (44)
  • While the saddle point is otherwise.
The conditions of stability for the fixed point $R_{5}$ are as follows:
  • For Stable node:
    $n\leq 0\,\,\text{and}\,\,\left(\alpha<-3\,\,\text{or}\,\,0<\alpha<-\frac{3n^{2}}{2n-1}\right),$ (45)
    $0<n<\frac{1}{2}\,\,\text{and}\,\,\alpha<-3,$ (46)
    $\frac{1}{2}<n<1\,\,\text{and}\,\,-\frac{3n^{2}}{2n-1}<\alpha<-3.$ (47)
  • For Unstable node:
    $\frac{1}{5}<n<\frac{1}{2}\,\,\text{and}\,\,\frac{1}{3}(8n-1)<\alpha<-\frac{3n^{2}}{2n-1},$ (48)
    $n>\frac{1}{2}\,\,\text{and}\,\,\alpha>\frac{1}{3}(8n-1).$ (49)
  • While the saddle point is otherwise.

### v2

@@
 Abstract
-In this work, we explore the behavior of interacting dark energy and dark matter within a model of $f(Q)$ gravity, employing a standard framework of dynamical system analysis.
-We consider the power-law $f(Q)$ model incorporating with two different forms of interacting dark energy and dark matter: $3\alpha H\rho_{m}$ and $\frac{\alpha}{3H}\rho_{m}\rho_{DE}$.
+In this work, we explore the behavior of interacting dark energy (DE) and dark matter (DM) within a model of $f(Q)$ gravity, employing a standard framework of dynamical system analysis.
+We consider the power-law $f(Q)$ model incorporating with two different forms of interacting DE and DM: $3\alpha H\rho_{m}$ and $\frac{\alpha}{3H}\rho_{m}\rho_{DE}$.
 The evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$ has been examined.
-Our results show that the universe was dominated by matter in the early stages and will be dominated by dark energy in later stages.
+Our results show that the universe was dominated by matter in the early stages and will be dominated by DE in later stages.
 Using the observational data, the fixed points are found to be stable and can be represented the de Sitter and quintessence acceleration solutions.
-We discover that the dynamical profiles of the universe in $f(Q)$ dark energy models are influenced by both the interaction term and the relevant model parameters.
+We discover that the dynamical profiles of the universe in $f(Q)$ DE models are influenced by both the interaction term and the relevant model parameters.
 Keywords: $f(Q)$ Gravity; Dark Energy; Dark Matter; Dynamical Analysis
-June 4, 2024
+January 3, 2025
 I Introduction
-Observational data has confirmed that the universe has entered a period of accelerated expansion.
-This evidence comes from various independent sources, including Supernovae Type Ia (SNe Ia) [1, 2, 3], temperature anisotropies in the Cosmic Microwave Background (CMB) observed by WMAP [4, 5], and Baryon Acoustic Oscillations [6, 7].
-According to the framework of general relativity (GR), this accelerated expansion is driven by a new energy density component with negative pressure, referred to as dark energy (DE).
-Despite its crucial role, little is known about DE.
-Understanding the origin of DE, which causes the current cosmic acceleration, remains one of the major unresolved issues in modern cosmology.
-The standard equation of state $\omega=p/\rho$, where $\omega\geq 0$, cannot explain this phase of cosmic acceleration.
-Instead, an unknown component with negative pressure, characterized by $\omega<-\tfrac{1}{3}$, is necessary to describe the universe’s acceleration.
-However, the exact nature of DE is still a mystery.
-Numerous attempts have been made to address this significant problem, as discussed in Refs.[8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
-The most straightforward candidate for dark energy is the cosmological constant, $\Lambda$.
-However, as pointed out in Ref.[23], the cosmological constant is vastly smaller than what is predicted by modern elementary particle theories.
-To differentiate between the cosmological constant and other potential sources of dark energy that change over time, researchers use dynamical models to study the evolution of the equation of state of DE.
-Therefore, a mechanism must be identified to achieve a small value of $\Lambda$ that aligns with observations.
-It is well established that the pure cosmological constant cannot explain the accelerated expansion of the early universe, as it cannot connect to the radiation-dominated era.
-An alternative and compelling candidate for both dark energy and inflation is a scalar field with a slowly varying potential.
-In the realm of dark energy, numerous scalar field models have been proposed, including quintessence [24, 25, 26, 27, 28, 29] and k-essence [30, 31, 32, 33] scenarios.
-These models predict various behaviors for the equation of state of dark energy.
-However, current observational data cannot distinguish these models from the $\Lambda$-cold-dark-matter ($\Lambda$CDM) model.
-Moreover, achieving viable scalar-field models within the framework of particle physics is challenging due to the extremely small scalar mass required for present-day cosmic acceleration [34, 35].
-Another approach to explaining the universe’s accelerated expansion involves dynamical dark energy models based on modifying gravity over large distances.
-Examples of such models include $f(R)$ gravity, where the Ricci scalar is replaced by a more general function, $f(R)$ [36, 37, 38, 39], as well as scalar-tensor theories [40, 41, 42, 43, 44, 45], Galileon gravity [46], and Gauss-Bonnet gravity [47, 48].
-For several decades, dynamical system analysis has been used in cosmology to qualitatively study these models, proving effective in identifying and classifying their asymptotic behaviors [49, 50].
-Recently, this framework has been applied to investigate various dark energy models, including those based on $f(R)$ gravity [13, 51].
-A series of papers [52, 54, 53, 55, 56, 57] has systematically analyzed $f(R)$ theories, identifying models that provide correct cosmological evolution from a vast number of possibilities.
-These analyses have led to the construction of cosmologically viable $f(R)$ gravity models, which feature appropriate trajectories in the dynamical system phase space and are compatible with local gravity constraints [20, 21, 55].
-A notable advantage of these models is that they achieve cosmic acceleration without requiring a dark energy component.
-Furthermore, these models impose stringent constraints from local gravity tests and various observational data, making them more rigorous compared to modified matter models.
-The authors of Refs.[58, 59, 60, 61, 62] initiated and expanded the phase space analysis of $f(R)$ gravity using the autonomous dynamical systems approach.
-While there has been considerable research on dynamical systems in $f(R)$ gravity involving non-interacting dark energy and dark matter [51], the study of interacting dark energy within the $f(R)$ gravity framework has received less attention, particularly in the context of dynamical system analysis [63, 64, 65, 66, 67, 68].
-Investigating interacting dark energy and dark matter is valuable for understanding the behavior of accelerated scaling solutions where the energy densities of dark energy ($\Omega_{DE}$) and dark matter ($\Omega_{DM}$) are approximately equal, and the universe is accelerating.
-This research could also elucidate the coupling between dark matter structure formation and the time evolution of dark energy, potentially explaining why dark energy dominates over dark matter in the late universe.
-Additionally, the impact of interacting dark energy models on the global 21cm signal was explored in Ref.[70], and some interacting dark energy models and their potential to cause future singularities have been analyzed in Ref.[71].
-The purpose of this work is to investigate the interaction between dark energy and dark matter within viable models of $f(Q)$ gravity.
-The $f(Q)$ gravity theory [72, 73] is one of the most promising approaches to explaining the accelerated expansion of the universe, involving dynamical dark energy models that modify gravity at large distances.
-This theory extends the alternative to General Relativity known as the symmetric teleparallel theory, which is based on the nonmetricity scalar $Q$, with both curvature and torsion being absent.
+The $\Lambda$-cold dark matter ($\Lambda$CDM) model has been extensively validated across a wide range of scales, from cosmic to very small, through various observations, including type Ia supernovae (SNe Ia) Riess et al.
+(1998, 1999); Perlmutter et al.
+(1999), baryon acoustic oscillations (BAO) Eisenstein et al.
+(2005); Percival et al.
+(2010), and the cosmic microwave background (CMB) Spergel et al.
+(2003); Komatsu et al.
+(2011).
+While $\Lambda$CDM has successfully explained numerous cosmological phenomena, it still encounters two significant challenges: (i) why does the cosmological constant observed today differ vastly from its theoretical prediction?
+and (ii) why are the current densities of dark matter (DM) and dark energy (DE) of the same magnitude?
+DM plays a crucial role in the formation of cosmic structures, while DE is responsible for driving the current accelerated expansion of the universe.
+While the nature of DM is partially understood through its indirect gravitational effects, DE remains profoundly enigmatic.
+Consequently, numerous cosmological models have been proposed and studied over recent years.
+The simplest approach involves non-interacting models, where DM and DE are conserved independently, resulting in separate evolutionary paths for these components.
+In contrast, more generalized models allow for interactions between DM and DE, providing a broader framework for understanding their dynamics.
+The interaction between DM and DE is a promising mechanism to address the cosmic coincidence problem, although its initial motivation was to resolve the discrepancy in the cosmological constant del Campo et al.
+(2009); Pavon and Zimdahl (2005).
+Early work by Wetterich demonstrated that an interaction between a scalar field and gravity could result in a dynamic effective cosmological constant that asymptotically approaches a small value, providing a plausible explanation for the mismatched cosmological constant Wetterich (1995).
+These dual motivations laid the foundation for exploring interactions within the dark sector.
+Initially, the primary motivation for interacting dark energy (IDE) models was to address or mitigate the coincidence problem.
+However, more recently, the focus has shifted toward resolving the discrepancy between the Hubble constant values derived from CMB and local measurements.
+As the tension between high-redshift and low-redshift Hubble constant measurements has persisted with improving data, IDE models have emerged as promising candidates to reconcile these discrepancies with the $\Lambda$CDM model Vagnozzi et al.
+(2017); Aljaf et al.
+(2021); Cheng et al.
+(2020); Yang et al.
+(2017); Nong and Liang (2024).
+The interaction within the dark sector is largely a phenomenological concept, as no fundamental principle explicitly dictates its existence.
+However, from a theoretical perspective, particularly in particle physics, any two matter fields, such as DM and DE fields, can interact.
+This idea has garnered significant attention within the cosmological community because of its potential implications.
+Notably, allowing such an interaction can transition the DE equation of state (EoS) from the quintessence regime to the phantom regime, effectively introducing a quintom-like behavior.
+Phenomenological models are typically formulated by incorporating an energy exchange between dark components into their continuity equations with an interacting kernel $\mathcal{U}$.
+Similarly to how interactions behave in particle physics, one would expect the kernel to be a function of the energy densities, involved $\rho_{DE}$, $\rho_{CDM}$ and of time, $H^{-1}$.
+Moreover, within the framework of field theory, considering interactions between the dark sectors is both natural and unavoidable.
+Exploring these interactions could provide valuable insights into the fundamental nature of DM and DE.
+Various interaction models that have been proposed and tested in the literature are linear models, such as $\mathcal{U}=\alpha_{m}H\rho_{m}$, $\mathcal{U}=\alpha_{DE}H\rho_{DE}$ and $\mathcal{U}=H(\alpha_{m}\rho_{m}+\alpha_{DE}\rho_{DE})$ Amendola et al.
+(2007); Quartin et al.
+(2008); Pavon and Wang (2009); Wang et al.
+(2016).
+Although only a few non-linear models have been proposed and investigated in the literature Wang et al.
+(2024).
+Another approach to explaining the accelerated expansion of the universe involves dynamical dark-energy models based on modifying gravity over large distances.
+Examples of such models include $f(R)$ gravity Santos et al.
+(2007); Sotiriou (2006); Capozziello et al.
+(2007, 2008), where the Ricci scalar is replaced by a more general function of $R$ as well as scalar-tensor theories Amendola (1999); Bartolo and Pietroni (2000); Nojiri et al.
+(2015); Baker and Harrison (2021), Galileon gravity Nicolis et al.
+(2009), and Gauss-Bonnet gravity Nojiri et al.
+(2005); Nojiri and Odintsov (2005).
+There are other different modified theories of gravity such as $f(R,T)$ gravity Harko et al.
+(2011); Alvarenga et al.
+(2013); Shabani and Farhoudi (2014); Fisher and Carlson (2019); Odintsov and Sáez-Gómez (2013); Barrientos O.
+and Rubilar (2014), $f(\mathcal{T})$ gravity Cai et al.
+(2016); Paliathanasis et al.
+(2016); Li et al.
+(2011); Capozziello et al.
+(2011), where $\mathcal{T}$ is a torsion, $f(Q)$ gravity Heisenberg (2024); Jimenez et al.
+(2018); Mandal et al.
+(2020); Gadbail et al.
+(2024); Lazkoz et al.
+(2019); Frusciante (2021); Sokoliuk et al.
+(2023), where $Q$ is the nonmetricity.
+For several decades, dynamical system analysis has been used in cosmology to qualitatively study these models, proving to be effective in identifying and classifying their asymptotic behaviors Wainwright and Lim (2005); Coley (1999).
+Studying autonomous dynamical analysis and point stability is essential for understanding the evolution of the universe and the behavior of cosmological models.
+Autonomous dynamical systems reformulate the equations governing cosmic evolution into first-order differential equations, simplifying the analysis and enabling a deeper exploration of the universe’s dynamics.
+Critical points within these systems represent key asymptotic behaviors of the universe, such as matter domination, radiation domination, or accelerated expansion.
+Stability analysis of these points is critical for determining whether the universe can evolve toward specific states, such as the current dark-energy-dominated epoch, thereby linking theoretical predictions with observational data.
+Moreover, dynamical system analysis provides insights into the influence of model parameters, such as coupling constants or power law indices, on the universe’s trajectory, helping identify ranges that lead to physically viable solutions, including late-time acceleration or scaling behaviors.
+This framework has been applied to investigate various DE models in various modified theories of gravity Beltrán Jiménez et al.
+(2020); Bahamonde et al.
+(2018); Samart et al.
+(2021); Khyllep et al.
+(2022); Hussain et al.
+(2024); An et al.
+(2016); Carloni et al.
+(2016); D’Ambrosio et al.
+(2022); Arora et al.
+(2022).
+The purpose of this work is to investigate the interaction between DE and DM within viable models of $f(Q)$ gravity Beltrán Jiménez et al.
+(2018, 2020).
+This theory extends the alternative to General Relativity (GR) known as the symmetric teleparallel theory, which is based on the nonmetricity scalar $Q$, with both curvature and torsion absent.
 Geometrically, the nonmetricity $Q$ describes the variation in the length of a vector during parallel transport.
-$f(Q)$ gravity presents intriguing applications [74, 75, 76, 77, 78, 79] and easily satisfies the constraints from gravitational wave observations [80].
-At cosmological scales, $f(Q)$ can create a gravitational interaction that reveals new and interesting patterns in both the background evolution [72, 73, 81, 82, 83, 84] and the propagation of scalar and tensor perturbations [73, 81, 74, 85, 86, 87].
-Additionally, certain forms of the $f(Q)$ function have been shown to mitigate the $\sigma_{8}$ tension [88], while others provide a better fit to cosmological data [89, 90, 91, 92, 93, 94, 95, 96, 97].
-This theory is thus challenging $\Lambda$CDM from different perspectives [98].
-Furthermore, a dynamical system analysis of both background and perturbation equations is conducted in $f(Q)$ gravity to independently verify the validity of result [87, 99, 100, 101, 102, 103, 104].
-The structure of this work is organized as follows: In Section II, we present the fundamental cosmological equations of the general $f(Q)$ theory and derive the Friedmann equation corresponding to the FLRW metric.
-In Section III, we derive the autonomous dynamical system within the framework of $f(Q)$ gravity theory, focusing on the interaction between dark energy and dark matter.
+Several studies have been done and explored in $f(Q)$ gravity which presents intriguing applications.
+For review, one can check the references Harko et al.
+(2018); Ayuso et al.
+(2021); Anagnostopoulos et al.
+(2021); Gadbail et al.
+(2022); Khyllep et al.
+(2023); Gadbail et al.
+(2023); Arora and Sahoo (2022); Barros et al.
+(2020); Atayde and Frusciante (2023); Nunes (2018).
+The structure of this work is organized as follows: In Section II, we present the fundamental cosmological equations of the general $f(Q)$ theory and derive the Friedmann equations corresponding to the FLRW metric.
+In Section III, we derive the autonomous dynamical system within the framework of $f(Q)$ gravity theory, focusing on the interaction between DE and DM.
 In Section IV, we apply the power-law $f(Q)$ model to enclose the dynamical system and conduct a further study of $f(Q)$ gravity.
 Finally, we discuss our findings in Section V.
-II Formulation of $f(Q)$ gravity theory
+II Formulation of $f(Q)$ gravity and FLRW cosmology
 In this section, we briefly discuss the formulation of $f(Q)$ gravity, a simple generalization of symmetric teleparallel gravity theory.
 This theory requires a different curvature and torsion-free connection, i.e., it is wholly dependent on nonmetricity.
-This formulation introduces new dynamics compared to General Relativity (GR) by allowing the function $f(Q)$ to dictate how nonmetricity influences the gravitational interaction.
-The non-metricity tensor $Q_{\sigma\mu\nu}$ is defined as
-$Q_{\sigma\mu\nu}=\nabla_{\sigma}g_{\mu\nu},$ (1)
-which geometrically describes the variation of the length of a vector in the parallel transport.
+This formulation introduces new dynamics compared to GR by allowing the function $f(Q)$ to dictate how nonmetricity influences the gravitational interaction.
+The non-metricity tensor $Q_{\sigma\mu\nu}$ is defined as $Q_{\sigma\mu\nu}=\nabla_{\sigma}g_{\mu\nu}$, which geometrically describes the variation of the length of a vector in the parallel transport.
+We begin by recalling that the general affine connection allows for a decomposition Nester and Yo (1999), which can be systematically expressed as follows:
+$\hat{\Gamma}^{\,\sigma}_{\,\,\,\mu\nu}=\Gamma^{\,\sigma}_{\,\,\,\mu\nu}+K^{\,\sigma}_{\,\,\,\mu\nu}+L^{\,\sigma}_{\,\,\,\mu\nu},$ (1)
+where the Levi-Civita connection $\Gamma^{\,\sigma}_{\,\,\,\mu\nu}$ is defined as
+$\Gamma^{\,\sigma}_{\,\,\,\mu\nu}=\frac{1}{2}g^{\sigma\lambda}\left(\partial_{\mu}g_{\lambda\nu}+\partial_{\nu}g_{\lambda\mu}-\partial_{\lambda}g_{\mu\nu}\right),$ (2)
+which can be uniquely determined by the first-order derivatives of the metric tensor $g_{\mu\nu}$.
+The contortion $K^{\,\sigma}_{\,\,\,\mu\nu}$ and deformation tensor $L^{\,\sigma}_{\,\,\,\mu\nu}$ are defined as
+$K^{\,\sigma}_{\,\,\,\mu\nu}$ $=$ $\frac{1}{2}T^{\,\sigma}_{\,\,\,\mu\nu}+T^{\,\,\,\,\,\,\,\sigma}_{(\mu\,\,\,\,\,\,\nu)},$
+$L^{\,\sigma}_{\,\,\,\mu\nu}$ $=$ $-\frac{1}{2}g^{\sigma\lambda}\left(Q_{\mu\lambda\nu}+Q_{\nu\lambda\mu}-Q_{\lambda\mu\nu}\right),$
+respectively, which describes non-Riemannian properties in the manifold.
+The contortion tensor disappears in the symmetric teleparallel theory because it follows an anti-symmetric property.
+The interplay between nonmetricity and the absence of torsion would influence cosmological models and the evolution of the universe.
+These effects could manifest in scenarios such as the dynamics of inflation, the behavior of DE, and the formation of large-scale structures.
 With the aid of the nonmetricity tensor, we can define the superpotential tensor $P_{\,\,\mu\nu}^{\sigma}$ as
 $4P_{\,\,\mu\nu}^{\sigma}=-Q^{\sigma}_{\,\,\,\,\mu\nu}+2Q^{\,\,\,\,\,\,\sigma}_{(\mu\,\,\,\,\nu)}-Q^{\sigma}g_{\mu\nu}-\tilde{Q}^{\sigma}g_{\mu\nu}-\delta^{\sigma}_{(\mu}\,Q\,_{\nu)},$ (3)
-where $Q_{\sigma}=Q_{\sigma\,\,\,\,\mu}^{\,\,\,\,\mu}$ and $\tilde{Q}_{\sigma}=Q^{\mu}_{\,\,\,\,\sigma\mu}\,$.
-From the above quantities, we can obtain a non-metricity scalar as
+where $Q_{\sigma}=Q_{\sigma\,\,\,\,\mu}^{\,\,\,\,\mu}$ and $\tilde{Q}_{\sigma}=Q^{\mu}_{\,\,\,\,\sigma\mu}\,$. From the above quantities, we can obtain a non-metricity scalar as
 $Q=-Q_{\sigma\mu\nu}P^{\sigma\mu\nu}.$ (4)
-The action for $f(Q)$ gravity [72]
-$S=\int\left[\frac{1}{2\kappa^{2}}f(Q)+\mathcal{L}_{m}\right]\sqrt{-g}\,d^{4}x,$ (4)
+The action for $f(Q)$ gravity Beltrán Jiménez et al. (2020)
+$S=\int\left[-\frac{1}{2\kappa^{2}}f(Q)+\mathcal{L}_{m}\right]\sqrt{-g}\,d^{4}x,$ (5)
 where $f(Q)$ represents any function of the scalar $Q$, $g$ denotes the determinant of $g_{\mu\nu}$, and $\mathcal{L}_{m}$ stands for the matter Lagrangian density.
-The equations of motion in $f(Q)$ gravity are derived by varying the action with respect to the metric and the connection.
-This leads to a set of modified field equations that incorporate the effects of nonmetricity, providing a richer structure for modeling gravitational phenomena, and it is written as
-$\frac{2}{\sqrt{-g}}\nabla_{\sigma}\left(f_{Q}\sqrt{-g}\,P^{\sigma}_{\,\,\mu\nu}\right)+\frac{1}{2}f(Q)\,g_{\mu\nu}\\ +f_{Q}\left(P_{\mu\sigma\lambda}Q_{\nu}^{\,\,\,\sigma\lambda}-2Q_{\sigma\lambda\mu}P^{\sigma\lambda}_{\,\,\,\,\,\,\nu}\right)=-T_{\mu\nu},$ (5)
+Here $\kappa^{2}=8\pi G$, where $G$ is the Newtonian gravitational constant.
+The equations of motion in $f(Q)$ gravity are derived by varying the action with respect to the metric.
+For simplicity, we set $18\pi G=$ 1.
+This leads to a set of modified field equations that incorporate the effects of non-metricity, providing a richer structure for modeling gravitational phenomena, and is written as
+$\frac{2}{\sqrt{-g}}\nabla_{\sigma}\left(f_{Q}\sqrt{-g}\,P^{\sigma}_{\,\,\mu\nu}\right)+\frac{1}{2}f(Q)\,g_{\mu\nu}+f_{Q}\left(P_{\mu\sigma\lambda}Q_{\nu}^{\,\,\,\sigma\lambda}-2Q_{\sigma\lambda\mu}P^{\sigma\lambda}_{\,\,\,\,\,\,\nu}\right)=T_{\mu\nu},$ (6)
 where $f_{Q}=\frac{df}{dQ}$. The energy-momentum tensor for matter is now defined as $T_{\mu\nu}\equiv-\frac{2}{\sqrt{-g}}\frac{\delta(\sqrt{-g})\mathcal{L}_{m}}{\delta g^{\mu\nu}}$.
+Also, by varying action (5) with respect to the connection results in
+$\nabla_{\sigma}\lambda_{k}^{\,\,\,\mu\nu\sigma}+\lambda_{k}^{\,\,\,\mu\nu}=\sqrt{-g}f_{Q}\,P_{\,k}^{\,\,\,\mu\nu}+H_{\,k}^{\,\,\,\mu\nu},$ (7)
+where $H_{\,k}^{\,\,\,\mu\nu}=-\frac{1}{2}\frac{\delta\mathcal{L}_{m}}{\delta\Gamma^{k}_{\,\,\,\mu\nu}}$ is the hypermomentum tensor density.
+It is possible to simplify Eq. (7) by taking into account the antisymmetry property of $\mu$ and $\nu$ in the Lagrangian multiplier coefficients
+$\nabla_{\mu}\nabla_{\nu}\left(f_{Q}\sqrt{-g}\,P_{\,k}^{\,\,\,\mu\nu}+H_{\,k}^{\,\,\,\mu\nu}\right)=0.$ (8)
+More specifically, the connection can be parameterized with a collection of functions $\xi^{\alpha}$ as $\hat{\Gamma}^{\,\sigma}_{\,\,\,\alpha\beta}=\frac{\partial x^{\sigma}}{\partial\xi^{\mu}}\partial_{\alpha}\partial_{\beta}\xi^{\mu}$.
+The connection equation of motion can be easily calculated by noticing that the variation of the connection with respect to $\xi^{\sigma}$ is equivalent to performing a diffeomorphism so that $\delta_{\xi}\hat{\Gamma}^{\,\sigma}_{\,\,\,\mu\nu}=-\mathcal{L}_{\xi}\hat{\Gamma}^{\,\sigma}_{\,\,\,\mu\nu}=-\nabla_{\mu}\nabla_{\nu}\xi^{\sigma}$, where we have used that the connection is flat and torsion-free Beltrán Jiménez et al.
+(2020).
+Furthermore, in the absence of hypermomentum, the connection field equations read as
+$\nabla_{\mu}\nabla_{\nu}\left(f_{Q}\sqrt{-g}\,P_{\,k}^{\,\,\,\mu\nu}\right)=0.$ (9)
+Symmetric teleparallel gravity provides a geometric formulation of gravity that is fully equivalent to general relativity.
+This equivalence becomes apparent in the coincident gauge, where the connection satisfies $\hat{\Gamma}^{\,\sigma}_{\,\,\,\mu\nu}=0$.
+Imposing the condition that the connection is symmetric eliminates the torsion tensor, allowing the Levi-Civita connection to be expressed in terms of the disformation tensor as $\hat{\Gamma}^{\,\sigma}_{\,\,\,\mu\nu}=-L^{\,\sigma}_{\,\,\,\mu\nu}$.
+Consequently, the non-metricity simplifies to $Q_{\sigma\mu\nu}=\partial_{\sigma}g_{\mu\nu}$.
+The use of the coincident gauge in $f(Q)$ gravity offers significant advantages, such as simplifying calculations by eliminating the connection, reducing ambiguities, and isolating the effects of non-metricity.
+This gauge allows for more tractable analytical and numerical analyses, making it particularly useful for cosmological studies.
+However, it comes with limitations, including a loss of generality, as results are gauge-dependent and may not fully capture the broader implications of the connection in $f(Q)$ gravity.
+Additionally, the coincident gauge may introduce biases and restrict the exploration of the connection’s physical role and the general behavior of the non-metricity scalar $Q$.
 To apply $f(Q)$ gravity in a cosmological context, we consider the spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) spacetime, characterized by the metric
@@
 where $a(t)$ is the cosmological scale factor.
-For this metric, the corresponding nonmetricity scalar is given by $Q=6H^{2}$, with $H=\frac{\dot{a}}{a}$ representing the Hubble parameter and the dot denotes a derivative with respect to the coordinate time $t$.
-Applying the FLRW metric into the general field equation (5), the Friedman equations of $f(Q)$ cosmology read as
-$6H^{2}f_{Q}-\frac{f}{2}$ $=$ $\rho$ (7)
+For this metric, the corresponding non-metricity scalar is given by $Q=6H^{2}$, with $H=\frac{\dot{a}}{a}$ representing the Hubble parameter, and the dot denotes a derivative with respect to the coordinate time $t$.
+Applying the FLRW metric to the general field equation (6), the Friedman equations of $f(Q)$ cosmology read as
+$6H^{2}f_{Q}-\frac{f}{2}$ $=$ $\rho\,,$ (11)
 $\left(12H^{2}\,f_{QQ}+f_{Q}\right)\dot{H}$ $=$ $-\frac{1}{2}(p+\rho),$ (12)
-where $f_{Q}=\frac{df}{dQ}$, and $f_{QQ}=\frac{d^{2}f}{dQ^{2}}$.
-To study the interaction between dark energy and dark matter, we have to modify the conservation equations of the dark energy and the matter given above by adding some coupling term $\mathcal{U}$ as [68]
+where $\rho$ and $p$ are energy density and pressure for the perfect fluid, respectively, while $f_{Q}=\frac{df}{dQ}$, and $f_{QQ}=\frac{d^{2}f}{dQ^{2}}$.
+To study the interaction between DE and DM, we have to modify the conservation equations of the DE and the DM given above by adding some coupling term $\mathcal{U}$ as Samart et al. (2021)
 $\dot{\rho}_{DE}+3H(p_{DE}+\rho_{DE})=\mathcal{U}\,,$
@@
 $\dot{\rho}_{r}+4H\,\rho_{r}=0\,.$
-In addition, the coupling term $\mathcal{U}$ between dark energy and dark matter can be interpreted as the exchange rate of energy density between these two components.
-When $\mathcal{U}>0$, energy is transferred from dark matter to dark energy, whereas when $\mathcal{U}<0$, energy is transferred from dark energy to dark matter.
-III Autonomous dynamical system of interacting dark energy and dark matter
-In this section, we derive the autonomous dynamical system within the framework of $f(Q)$ gravity theory, focusing on the interaction between dark energy and dark matter.
+In addition, the coupling term $\mathcal{U}$ between DE and DM can be interpreted as the exchange rate of energy density between these two components.
+When $\mathcal{U}>0$, energy is transferred from DM to DE, whereas when $\mathcal{U}<0$, energy is transferred from DE to DM.
+III Autonomous dynamical system of interacting DE and DM
+In this section, we derive the autonomous dynamical system within the framework of $f(Q)$ gravity theory, focusing on the interaction between DE and DM.
 Our approach involves incorporating two interacting terms.
 The first term relies solely on the energy density of the matter sector, the Hubble parameter, and a coupling constant $\alpha$, structured multiplicatively as $\mathcal{U}=3\alpha\,H\,\rho_{m}$.
-The second term encompasses the multiplication of energy densities from both sectors, capturing the immediate effects of both dark matter and dark energy on the interaction term, with a dimensionless coupling constant $\alpha$, denoted as $\mathcal{U}=\frac{\alpha}{3H}\rho_{m}\,\rho_{de}$.
+The second term encompasses the multiplication of energy densities from both sectors, capturing the immediate effects of both DM and DE on the interaction term, with a dimensionless coupling constant $\alpha$, denoted as $\mathcal{U}=\frac{\alpha}{3H}\rho_{m}\,\rho_{DE}$.
 According to the Friedmann equation (11), we can define the dimensionless variable as
-$x=\frac{f}{12H^{2}\,F},\,\,\,\,\,y=\frac{\rho_{r}}{6H^{2}\,F}.$ (9)
-Invoking the above dimensionless variables, it is straightforward to derive the autonomous equations which play a key role for studying the dynamical system for the interacting dark energy and dark matter.
+$x=\frac{f}{12H^{2}\,f_{Q}},\,\,\,\,\,y=\frac{\rho_{r}}{6H^{2}\,f_{Q}}.$ (13)
+Invoking the above dimensionless variables, it is straightforward to derive the autonomous equations which play a key role for studying the dynamical system for the interacting DE and DM.
 III.1 Case I : $\mathcal{U}=3\alpha\,H\,\rho_{m}$
@@
 $\frac{dy}{dN}=-\frac{\dot{H}}{H^{2}}\,y-y+y^{2}-3xy+\alpha\,xy(1-x-y).$ (17)
-where the variable $N=ln\,a$ is e-folding number and leads to $\frac{d}{dN}=\frac{1}{H}\frac{d}{dt}$.
 Furthermore, we obtain the generalized equation for $\frac{\dot{H}}{H^{2}}$ as
@@
 $\Omega_{m}=1-x-y,\,\,\,\,\,\,\,\Omega_{r}=y,\,\,\,\,\,\,\,\Omega_{DE}=x.$ (19)
-Finally, we define the equation of state and deceleration parameters corresponding to the dimensionless variables to check the appropriate acceleration expansion of the universe, that is
+Finally, we define the EoS and deceleration parameters corresponding to the dimensionless variables to check the appropriate acceleration expansion of the universe, that is
 $\omega=-1+\frac{2f_{Q}(1-x+\frac{y}{3})}{2Qf_{QQ}+f_{Q}}$ (20)
@@
 Both $\omega$ and $q$ represent different aspects of cosmic evolution, each uniquely influencing the large-scale structure of the universe.
-Grasping the significance of these parameters is essential for studying the effects of dark energy across various stages of cosmic evolution.
-Moreover, these parameters aid in comparing and differentiating between various dark energy models, each employing distinct mechanisms to drive cosmic acceleration.
+Grasping the significance of these parameters is essential for studying the effects of DE across various stages of cosmic evolution.
+Moreover, these parameters aid in comparing and differentiating between various DE models, each employing distinct mechanisms to drive cosmic acceleration.
 IV Model:- $f(Q)=6\gamma\,H_{0}^{2}\left(\frac{Q}{Q_{0}}\right)^{n}$
 In the previous section, we derived autonomous dynamical systems, and our primary objective is to investigate and analyze them.
-To accomplish this goal, we will focus on the designated power-law $f(Q)$ model, which has the form $f(Q)=6\gamma\,H_{0}^{2}\left(\frac{Q}{Q_{0}}\right)^{n}$, where $\gamma$ and $n$ are free model parameters.
+To accomplish this goal, we will focus on the designated power-law $f(Q)$ model, which has the form $f(Q)=6\gamma\,H_{0}^{2}\left(\frac{Q}{Q_{0}}\right)^{n}$, where $\gamma$, $n$ are free model parameters and $Q_{0}=6H_{0}^{2}$.
 Corresponding to our power-law $f(Q)$ model, Eq. (18) can be simplified to obtain
 $\frac{\dot{H}}{H^{2}}=-\frac{3(1-x+\frac{1}{3}y)}{2n-1}.$ (22)
-Now, we will find the critical points and corresponding eigenvalues for both the dynamical systems.
-Studying the critical points and assessing their stability is crucial for a comprehensive understanding of critical aspects in cosmic evolution driven by dark energy and dark matter interactions, as explored within the scope of this study.
-The properties of the dynamical system depend on the values of the constants $\alpha$ and $n$.
+We will now derive the critical points of the system by setting $dx/dN=dy/dN=0$ and determine the corresponding eigenvalues for both dynamical systems.
+The critical points, which represent the solutions of the dynamical system, provide an initial qualitative understanding of the phase space.
+As discussed below, these points can be classified based on their stability properties.
+In the absence of singularities or strange attractors, the trajectories of $x(N)$ and $y(N)$, which are generally obtained numerically, tend to evolve from unstable fixed points to stable fixed points, possibly passing through intermediate saddle points.
+Analyzing the critical points and evaluating their stability is essential for a thorough understanding of the critical aspects of cosmic evolution driven by interactions between DE and DM, as explored in this study.
+Furthermore, the properties of the dynamical system depend significantly on the values of the constants $\alpha$ and $n$.
 IV.1 For case I
-The fixed points $(x,y)$ of the general dynamical system are the following:
-$P_{1}:(1,0),\,\,\,\,P_{2}:\left(\frac{1}{2n},\frac{2n-1}{2n}\right),$ (19)
-$P_{3}:\left(\frac{1}{2n+\alpha(2n-1)},0\right).$ (20)
-Furthermore, the corresponding eigenvalues for each fixed point are given by
-$P_{1}:\{-4,-3(\alpha+1)\},\,\,\,\,P_{2}:\{4,1-3\alpha\},$ (21)
-$P_{3}:\{3(\alpha+1),3\alpha-1\}.$ (22)
-Let us now focus on the dynamics of each critical point and its features in the further subsections.
+The fixed points $(x,y)$ of the general dynamical system are described in table 1. Let us now focus on the dynamics of each critical point and its features in the following subsections.
+Critical Points $\Omega_{m}$ $\Omega_{DE}$ $w$ Eigenvalues Stability
+$P_{1}:(1,0)$ 00 $1$ $-1$ $\{-4,\,\,\,-3(\alpha+1)\}$
+Stable node for $\alpha>-1$,
+Saddle for $\alpha<-1$,
+Non-hyperbolic for $\alpha=-1$.
+$P_{2}:\left(\frac{1}{2n},\frac{2n-1}{2n}\right)$ 00 $\frac{1}{2n}$ $-1+\frac{4}{3n}$$\{4,1-3\alpha\}$
+Saddle node for $\alpha>1/3$,
+Unstable for $\alpha<1/3$,
+Non-hyperbolic for $\alpha=1/3$.
+$P_{3}:\left(\frac{1}{2n+\alpha(2n-1)},0\right)$ $\frac{1}{\alpha-2(\alpha+1)n}+1$ $\frac{1}{2n+\alpha(2n-1)}$ $\frac{3\alpha-2(\alpha+1)n+2}{2(\alpha+1)n-\alpha}$$\{3(\alpha+1),3\alpha-1\}$
+Stable node for $\alpha<-1$,
+Saddle for $-1<\alpha<1/3$,
+Unstable node for $\alpha>1/3$.
+Table 1: This table summarizes the analysis of the critical points for Case I.
 IV.1.1 $P_{1}$: the de Sitter fixed point
@@
 $P_{1}:(x,y)=(1,0).$
-This critical point is independent of both the model and coupling parameters. Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters are
+This critical point is independent of both the model and coupling parameters. Corresponding to this fixed point, the matter density, DE density, and effective EoS parameters are
 $\Omega_{m}=0,\,\,\,\,\Omega_{DE}=1,\,\,\,\,\omega=-1.$ (23)
-At this point, both dark matter and radiation are absent, and the universe is dominated by dark energy.
-The equation of state (EoS) at this fixed point indicates that the universe is experiencing an accelerated expansion.
+At this point, both DM and radiation are absent, and the universe is dominated by DE.
+The EoS at this fixed point indicates that the universe is experiencing an accelerated expansion.
 To study the stability of this fixed point, we can analyze the eigenvalues of the Jacobian matrix, which describe the behavior of the fixed point.
-The eigenvalues for the fixed point $P_{1}$ read
-$\{-4,\,\,\,-3(\alpha+1)\}.$ (24)
-The stability behavior of this fixed point (which depends on $\alpha$) is obtained as:
+The eigenvalues for the fixed point $P_{1}$ read $\{-4,\,\,\,-3(\alpha+1)\}$.
+The stability behavior of this fixed point (which depends on $\alpha$) is obtained as
   • Stable node for $\alpha>-1$,
@@
 $P_{2}:(x,y)=\left(\frac{1}{2n},\frac{2n-1}{2n}\right).$
-The characteristics of this critical point solely depend on the model parameter $n$.
-Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters are obtained as
+The characteristics of this critical point solely depend on the model parameter $n$. Corresponding to this fixed point, the matter density, DE density, and effective EoS parameters are obtained as
 $\Omega_{m}=0,\,\,\,\,\Omega_{DE}=\frac{1}{2n},\,\,\,\,\omega=-1+\frac{4}{3n}.$ (24)
@@
 In this fixed-point solution, a matter-dominated era occurs for $n=4/3$ and a radiation-dominated era for $n=1$.
-The eigenvalues for the fixed point $P_{2}$ are as follows:
-{4,   $\{4,\,\,\,1-3\alpha\}.$ (26)
+The eigenvalues for the fixed point $P_{2}$ are as follows {4,   $\{4,\,\,\,1-3\alpha\}$.
 The stability behavior for this particular fixed point can be obtained for different $\alpha$, given by
@@
 IV.1.3 $P_{3}$: scaling solution fixed point
-It is worth noting that our result modifies the scaling solution of $f(Q)$ gravity with the interacting dark energy parameter $\alpha$. The fixed point $P_{3}$ reads
+It is worth noting that our result modifies the scaling solution of $f(Q)$ gravity with the interacting DE parameter $\alpha$. The fixed point $P_{3}$ reads
 $P_{3}:(x,y)=\left(\frac{1}{2n+\alpha(2n-1)},0\right).$
 The properties of this critical point are influenced by both the model parameter $n$ and the coupling parameter $\alpha$.
-Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters read as
-$\Omega_{m}=\frac{1}{\alpha-2(\alpha+1)n}+1,\,\,\,\,\Omega_{DE}=\frac{1}{2n+\alpha(2n-1)},$
-and
-$\omega=\frac{3\alpha-2(\alpha+1)n+2}{2(\alpha+1)n-\alpha}.$
+Corresponding to this fixed point, the matter density, DE density, and effective EoS parameters read as
+$\Omega_{m}=\frac{1}{\alpha-2(\alpha+1)n}+1,\quad\Omega_{DE}=\frac{1}{2n+\alpha(2n-1)},\quad\omega=\frac{3\alpha-2(\alpha+1)n+2}{2(\alpha+1)n-\alpha}.$
 The conditions for an accelerating universe are shown in Fig. 1.
-Figure 1: In this figure, we show the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe.
+Figure 1: The plot illustrates the relationship between the model parameter $n$ and the interaction parameter $\alpha$ for an accelerating universe.
 The light blue region represents the quintessence-like behavior of an accelerating universe (i.e., $-1<\omega<-1/3$), while the golden region represents the phantom-like behavior of an accelerating universe (i.e., $\omega<-1$).
-In this fixed point solution, we obtained a matter-dominated era for $n=\frac{3\alpha+2}{2(\alpha+1)}$, $1+\alpha\neq 0$, and radiation-dominated era for $n=\frac{5\alpha+3}{4(\alpha+1)}$, $1+\alpha\neq 0$.
-The eigenvalues for the fixed point $P_{3}$ read as
 […]
-We determine the stability of this fixed point for various ranges of $\alpha$. The ranges are as follows:
-  • Stable node for $\alpha<-1$
-  • Saddle for $-1<\alpha<1/3$
+  • Stable node for $\alpha<-1$,
+  • Saddle for $-1<\alpha<1/3$,
   • Unstable node for $\alpha>1/3$.
@@
 This is done using the numerical solution of the dynamical system with the initial conditions $x(0)=0.7$ and $y(0)=0.00005$.
-The motivation behind the choice of initial conditions is as follows: $x$ represents the dark energy density, which currently has a value of $0.7$.
+The motivation behind the choice of initial conditions is as follows: $x$ represents the DE density, which currently has a value of $0.7$.
 Similarly, $y$ represents the radiation density, which currently has a value of $0.00005$.
-Figure 2: The plot on the left displays density parameters for matter ($\Omega_{m}$), dark energy ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$.
-The right panel shows the deceleration parameter $q$ and the equation of state parameter $\omega$ as functions of redshift $z$.
-These two panels illustrate the trajectory for the positive coupling parameter, which can transfer energy from dark matter to dark energy.
-Figure 3: The left panel displays density parameters for matter ($\Omega_{m}$), dark energy ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$.
-The right panel shows the deceleration parameter $q$ and the equation of state parameter $\omega$ as functions of redshift $z$.
-These two panels illustrate the trajectory for the negative coupling parameter, which can transfer energy from dark energy to dark matter.
+Figure 2: The left panel shows the density parameters for matter ($\Omega_{m}$), DE ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift ($z$).
+The right panel displays the deceleration parameter ($q$) and the equation of state parameter ($\omega$) as functions of redshift ($z$).
+Together, these two panels illustrate the trajectory for a positive coupling parameter, which can facilitate the transfer of energy from DM to DE.
+Figure 3: The left panel displays density parameters for matter ($\Omega_{m}$), DE ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$.
+The right panel shows the deceleration parameter $q$ and the EoS parameter $\omega$ as functions of redshift $z$.
+These two panels illustrate the trajectory for the negative coupling parameter, which can transfer energy from DE to DM.
 For $n=3/2$ and $\alpha=1/2$ in Fig.
-2, $\alpha>0$ indicates that the coupling term $\mathcal{U}>0$, signifying energy transfer from dark matter to dark energy.
-This figure shows that the universe was dominated by matter in the early stages and will be dominated by dark energy in later stages.
-Currently, the universe is dominated by dark energy, with parameter values $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, and $\Omega_{DE}=0.7$, $q_{0}=-0.55$, and $\omega_{0}=-0.70$.
+2, $\alpha>0$ indicates that the coupling term $\mathcal{U}>0$, signifying energy transfer from DM to DE.
+This figure shows that the universe was dominated by matter in the early stages and will be dominated by DE in later stages.
+Currently, the universe is dominated by DE, with parameters value $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, $\Omega_{DE}=0.7$, $q_{0}=-0.55$, and $\omega_{0}=-0.70$.
 For these values, fixed point $P_{1}$ is stable and represents the de Sitter acceleration solution, while fixed point $P_{2}$ is a saddle-node, and $P_{3}$ is an unstable node that cannot demonstrate universal acceleration.
 In Fig.
-3, with $n=3/2$ and $\alpha=-2$, $\alpha<0$ indicates that the coupling term $\mathcal{U}<0$, that means energy transfers from dark energy to dark matter.
-This figure shows that the universe was dominated by dark energy in the early stages and will be dominated by dark matter at later stages.
-Currently, the universe remains dominated by dark energy, with parameter values $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, and $\Omega_{DE}=0.7$, $q_{0}=-0.563$, and $\omega_{0}=-0.71$.
+3, with $n=3/2$ and $\alpha=-2$, $\alpha<0$ indicates that the coupling term $\mathcal{U}<0$, that means energy transfers from DE to DM.
+This figure shows that the universe was dominated by DE in the early stages and will be dominated by DM at later stages.
+Currently, the universe remains dominated by DE, with parameters value $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, $\Omega_{DE}=0.7$, $q_{0}=-0.563$, and $\omega_{0}=-0.71$.
 Here, fixed point $P_{3}$ is stable and exhibits acceleration for the early and present universe but fails to show acceleration for late times, while fixed point $P_{1}$ is a saddle-node, and $P_{2}$ is an unstable node.
@@
 4 shows the phase space trajectories of the fixed points $P_{1}$, $P_{2}$, and $P_{3}$.
-Figure 4: Phase plot of $x$ vs $y$. The upper phase plot shows the stable behavior of $P_{1}$ for $\alpha>0$, and the lower phase plot shows the stable behavior of $P_{3}$ for $\alpha<-1$.
+Figure 4: Phase space of $x$ vs $y$.
+The left phase plot illustrates the stable behavior of $P_{1}$ when $\alpha>0$, while the right phase plot showcases the stable behavior of $P_{3}$ for $\alpha<-1$.
 IV.2 For case II
-Next, the second case in $f(Q)$ gravity is also investigated in detail for the interacting dark energy system.
-The models have been systematically constructed with possible ranges of their model parameters derived from the dynamical system analysis within the interacting dark energy framework.
+Next, the second case in $f(Q)$ gravity is also investigated in detail for the interacting DE system.
+The models have been systematically constructed with possible ranges of their model parameters derived from the dynamical system analysis within the interacting DE framework.
 It is important to examine how the interaction affects the cosmological viability of the $f(Q)$ gravity models.
-The fixed points $(x,y)$ of the dynamical system are the following:
-$R_{1}:(1,0),\,\,\,\,R_{2}:\left(\frac{1}{2n},\frac{2n-1}{2n}\right),R_{3}:\left(\frac{1}{\alpha},\frac{-3\alpha+8n-1}{\alpha}\right),$
-$R_{4}:\left(\frac{-\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{2\alpha n-\alpha},0\right),$ (28)
-$R_{5}:\left(\frac{\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{2\alpha n-\alpha},0\right).$
-The corresponding eigenvalues are
-$R_{1}:\{-4,-3-\alpha\},\,\,\,\,R_{2}:\{4,1-\frac{\alpha}{2n}\},$
-$R_{3}:\{\mathcal{A}_{1}-\frac{4n}{\alpha}+4,-\mathcal{A}_{1}-\frac{4n}{\alpha}+4\},$
-$R_{4}:\left\{\frac{5n-1+\mathcal{A}_{2}}{1-2n},\frac{-2\mathcal{A}_{2}^{2}-2(3n-\alpha+2n\alpha)\mathcal{A}_{2}}{(2n-1)^{2}\alpha}\right\},$
-$R_{5}:\left\{\frac{-2\mathcal{A}_{2}^{2}+2(3n-\alpha+2n\alpha)\mathcal{A}_{2}}{(2n-1)^{2}\alpha},\frac{1-5n+\mathcal{A}_{2}}{2n-1}\right\}.$
-where $\mathcal{A}_{1}=\frac{2\sqrt{(2n-1)\left(3(\alpha-1)\alpha^{2}+8n^{3}-4n^{2}-6(\alpha-1)\alpha n\right)}}{\alpha-2\alpha n}$ and $\mathcal{A}_{2}=\sqrt{9n^{2}-3\alpha+6n\alpha}$.
+The fixed points $(x,y)$ of the dynamical system are described in table 2.
+Critical Points $\Omega_{m}$ $\Omega_{DE}$ $w$ Eigenvalues Stability
+$R_{1}:(1,0)$ 00 $1$ $-1$ $\{-4,-3-\alpha\}$
+Stable node for $\alpha>-3$,
+Saddle for $\alpha<-3$,
+Non-hyperbolic for $\alpha=-3$.
+$R_{2}:\left(\frac{1}{2n},\frac{2n-1}{2n}\right)$ 00 $\frac{1}{2n}$ $-1+\frac{4}{3n}$$\{4,1-\frac{\alpha}{2n}\}$
+Saddle node for $\alpha>2n$,
+Unstable for $\alpha<2n$,
+Non-hyperbolic for $\alpha=2n$.
+$R_{3}:\left(\frac{1}{\alpha},\frac{-3\alpha+8n-1}{\alpha}\right)$ $4-\frac{8n}{\alpha}$ $\frac{1}{\alpha}$ $\frac{8}{3\alpha}-1$$\{\mathcal{A}_{1}-\frac{4n}{\alpha}+4,-\mathcal{A}_{1}-\frac{4n}{\alpha}+4\}$
+The stability conditions
+are shown in Fig. 5
+$R_{4}:\left(\frac{-\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{2\alpha n-\alpha},0\right)$ Sec.IV.2.4 IV.2.4 IV.2.4 $\left\{\frac{5n-1+\mathcal{A}_{2}}{1-2n},\frac{-2\mathcal{A}_{2}^{2}-2(3n-\alpha+2n\alpha)\mathcal{A}_{2}}{(2n-1)^{2}\alpha}\right\}$
+The stability conditions
+are shown in Fig. 5
+$R_{5}:\left(\frac{\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{2\alpha n-\alpha},0\right)$ Sec.IV.2.5 IV.2.5 IV.2.5 $\left\{\frac{-2\mathcal{A}_{2}^{2}+2(3n-\alpha+2n\alpha)\mathcal{A}_{2}}{(2n-1)^{2}\alpha},\frac{1-5n+\mathcal{A}_{2}}{2n-1}\right\}$
+The stability conditions
+are shown in Fig. 5
+Table 2: This figure summarizes the results of the analysis of the critical points for Case II.
+The critical points $R_{4}$ and $R_{5}$, however, are difficult to present comprehensively in a table format.
+Detailed analyses of these points can be found directly in Sections IV.2.4 and IV.2.5.
+The expressions of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ in the table are given by $\mathcal{A}_{1}=\frac{2\sqrt{(2n-1)\left(3(\alpha-1)\alpha^{2}+8n^{3}-4n^{2}-6(\alpha-1)\alpha n\right)}}{\alpha-2\alpha n}$ and $\mathcal{A}_{2}=\sqrt{9n^{2}-3\alpha+6n\alpha}$.
 IV.2.1 $R_{1}$: the de Sitter fixed point
-Corresponding to fixed point $R_{1}:(x,y)=(1,0)$, the matter density, dark energy density, and effective EoS parameters are
+Corresponding to fixed point $R_{1}:(x,y)=(1,0)$, the matter density, DE density, and effective EoS parameters are
 $\Omega_{m}=0,\,\,\,\,\Omega_{DE}=1,\,\,\,\,\omega=-1.$ (25)
-At this point, both dark matter and radiation are absent, and the universe is dominated by dark energy.
+At this point, both DM and radiation are absent, and the universe is dominated by DE.
 The EoS at this fixed point indicates that the universe is experiencing an accelerated expansion.
 To study the stability of this fixed point, we can analyze the eigenvalues of the Jacobian matrix, which describe the behavior of the fixed point.
-The eigenvalues for the fixed point $R_{1}$ read
-$\{{-4,\,\,\,-\alpha-3\}}.$ (30)
+The eigenvalues for the fixed point $R_{1}$ are $\{{-4,\,\,\,-\alpha-3\}}$.
 The stability behavior of this fixed point is as follows:
@@
 $R_{2}:(x,y)=\left(\frac{1}{2n},\frac{2n-1}{2n}\right)$
-The characteristics of this critical point solely depend on the model parameter $n$.
-Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters are obtained as
+The characteristics of this critical point solely depend on the model parameter $n$. Corresponding to this fixed point, the matter density, DE density, and effective EoS parameters are obtained as
 $\Omega_{m}=0,\,\,\,\,\Omega_{DE}=\frac{1}{2n},\,\,\,\,\omega=-1+\frac{4}{3n}.$ (26)
@@
 In this fixed-point solution, we have obtained a matter-dominated era for $n=4/3$ and a radiation-dominated era for $n=1$.
-The corresponding eigenvalues for the fixed point $R_{2}$ are as follows:
-{4,  $\{{4,\,\,1-\frac{\alpha}{2n}\}}.$ (32)
-We determine the stability of this fixed point for various ranges of $\alpha$. The ranges are as follows:
+The corresponding eigenvalues for the fixed point $R_{2}$ are {4,  $\{{4,\,\,1-\frac{\alpha}{2n}\}}$.
+We determine the stability of this fixed point for various ranges of $\alpha$.
+The ranges are as follows
   • Saddle node for $\alpha>2n$,
@@
 $R_{3}:(x,y)=\left(\frac{1}{\alpha},\frac{-3\alpha+8n-1}{\alpha}\right).$
-Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters read as
+Corresponding to this fixed point, the matter density, DE density, and effective EoS parameters read as
 $\Omega_{m}=4-\frac{8n}{\alpha},\,\,\,\,\Omega_{DE}=\frac{1}{\alpha},\,\,\,\,\omega=\frac{8}{3\alpha}-1.$ (27)
@@
 The fixed point $R_{4}$ represents the scaling solution point of the universe.
-In addition, it is worth noting that our result modifies the scaling solution with the interacting dark energy parameter $\alpha$.
+In addition, it is worth noting that our result modifies the scaling solution with the interacting DE parameter $\alpha$.
 The point $R_{4}$ is
 $R_{4}:(x,y)=\left(\frac{-\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{2\alpha n-\alpha},0\right).$
-Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters are obtained as
+Corresponding to this fixed point, the matter density, DE density, and effective EoS parameters are obtained as
 $\Omega_{m}=\frac{-\alpha+\sqrt{-3\alpha+9n^{2}+6\alpha n}+(2\alpha+3)n}{\alpha(2n-1)}\,,$
@@
 The conditions for an accelerating universe and the stability are shown in Figs.
-6 and 7.
-Figure 6: This plot shows the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe.
+6(a) and 6(b).
+Figure 6: The region plot (a) shows the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe.
 The light blue region represents the quintessence-like behavior of an accelerating universe (i.e., $-1<\omega<-1/3$), while the golden region represents the phantom-like behavior of an accelerating universe (i.e., $\omega<-1$).
-Figure 7: The region plot illustrates the stable, unstable, and saddle behaviors associated with the fixed point $R_{4}$.
+The region plot (b) illustrates the stable, unstable, and saddle behaviors associated with the fixed point $R_{4}$.
 IV.2.5 $R_{5}$: Additional scaling solution point
@@
 $R_{5}:(x,y)=\left(\frac{\sqrt{3}\sqrt{-\alpha+3n^{2}+2\alpha n}-3n}{2\alpha n-\alpha},0\right)$
-the matter density, dark energy density, and effective EoS parameters are given by
+the matter density, DE density, and effective EoS parameters are given by
 $\Omega_{m}=\frac{3n-\sqrt{-3\alpha+9n^{2}+6\alpha n}}{\alpha(2n-1)}+1\,,$
@@
 $w=\frac{-3\alpha-4\alpha n^{2}-2\sqrt{-3\alpha+9n^{2}+6\alpha n}+(8\alpha+6)n}{\alpha(1-2n)^{2}}\,.$
-The conditions for an accelerating universe and the stability are depicted in Figs. 8 and 9.
-Figure 8: The plot show the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe.
+The conditions for an accelerating universe and the stability are depicted in Figs. 7(a) and 7(b).
+Figure 7: The region plot (a) shows the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe.
 The light blue region represents the quintessence-like behavior of an accelerating universe (i.e., $-1<\omega<-1/3$), while the golden region represents the phantom-like behavior of an accelerating universe (i.e., $\omega<-1$).
-Figure 9: The region plot illustrates the stable, unstable, and saddle behaviors associated with the fixed point $R_{5}$.
-In Figures 10 and 11, we illustrate the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$.
-For $n=3/2$ and $\alpha=4$ in Figure 10, $\alpha>0$ indicates that the coupling term $\mathcal{U}>0$, signifying energy transfer from dark matter to dark energy.
-This figure shows that the universe was dominated by matter in the early stages and will be dominated by dark energy in later stages.
-Currently, the universe is dominated by dark energy, with parameter values $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, and $\Omega_{DE}=0.7$, $q_{0}=-0.548$, and $\omega_{0}=-0.696$.
+The region plot (b) illustrates the stable, unstable, and saddle behaviors associated with the fixed point $R_{5}$.
+In Figures 8 and 9, we illustrate the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$.
+For $n=3/2$ and $\alpha=4$ in Figure 8, $\alpha>0$ indicates that the coupling term $\mathcal{U}>0$, signifying energy transfer from DM to DE.
+This figure shows that the universe was dominated by matter in the early stages and will be dominated by DE in later stages.
+Currently, the universe is dominated by DE, with parameters value $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, $\Omega_{DE}=0.7$, $q_{0}=-0.548$, and $\omega_{0}=-0.696$.
 For these values, the fixed points $R_{1}$ and $R_{4}$ are stable and represent the de Sitter and quintessence acceleration solutions, respectively.
-In Figure 11, with $n=3/2$ and $\alpha=-4$, $\alpha<0$ indicates that the coupling term $\mathcal{U}<0$, meaning energy transfers from dark energy to dark matter.
-This figure shows that the universe was dominated by dark energy in the early stages and will be dominated by dark matter in later stages.
-Currently, the universe remains dominated by dark energy, with parameter values $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, and $\Omega_{DE}=0.7$, $q_{0}=-0.55$, and $\omega_{0}=-0.70$.
+In Figure 9, with $n=3/2$ and $\alpha=-4$, $\alpha<0$ indicates that the coupling term $\mathcal{U}<0$, meaning energy transfers from DE to DM.
+This figure shows that the universe was dominated by DE in the early stages and will be dominated by DM in later stages.
+Currently, the universe remains dominated by DE, with parameters value $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, $\Omega_{DE}=0.7$, $q_{0}=-0.55$, and $\omega_{0}=-0.70$.
 For these values, the fixed points $R_{4}$ and $R_{5}$ have imaginary values, and the remaining point cannot exhibit stable behavior.
-Figure 10: The upper panel displays density parameters for matter ($\Omega_{m}$), dark energy ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$.
-The lower panel shows the deceleration parameter $q$ and the equation of state parameter $\omega$ as functions of redshift $z$.
-These two panels illustrate the trajectory for the positive coupling parameter, which can transfer energy from dark matter to dark energy.
-Figure 11: The left panel displays density parameters for matter ($\Omega_{m}$), dark energy ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$.
-The right panel shows the deceleration parameter $q$ and the equation of state parameter $\omega$ as functions of redshift $z$.
-These two panels illustrate the trajectory for the negative coupling parameter, which can transfer energy from dark energy to dark matter.
-Figure 12: Phase plot of $x$ vs $y$. The phase plot above shows the stable behavior of $R_{1}$ and $R_{4}$ for $\alpha>0$.
+Figure 8: The left panel displays density parameters for matter ($\Omega_{m}$), DE ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$.
+The right panel shows the deceleration parameter $q$ and the EoS parameter $\omega$ as functions of redshift $z$.
+These two panels illustrate the trajectory for the positive coupling parameter, which can transfer energy from DM to DE.
+Figure 9: The left panel displays density parameters for matter ($\Omega_{m}$), DE ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$.
+The right panel shows the deceleration parameter $q$ and the EoS parameter $\omega$ as functions of redshift $z$.
+These two panels illustrate the trajectory for the negative coupling parameter, which can transfer energy from DE to DM.
+Figure 10: Phase plot of $x$ versus $y$. These phase plots depict the stable behavior of $R_{1}$ and $R_{4}$ for $\alpha>0$.
 V Conclusion
-In this work, we have explored the behavior of interacting dark energy and dark matter within $f(Q)$ gravity, employing a standard framework of dynamical system analysis.
-We have considered the power-law $f(Q)=6\gamma\,H_{0}^{2}\left(\tfrac{Q}{Q_{0}}\right)^{n}$ model incorporating with two different forms of interacting dark energy and dark matter: $3\alpha H\rho_{m}$ and $\frac{\alpha}{3H}\rho_{m}\rho_{DE}$.
-The parameter $\alpha$ in the interacting terms plays a crucial role in determining the viable conditions and estimating the transition from the matter-dominated era (saddle point) to the dark energy-dominated era (stable node) in viable gravity models at late times.
+In this work, we have explored the behavior of interacting DE and DM within $f(Q)$ gravity, employing a standard framework of dynamical system analysis.
+We have considered the power-law $f(Q)=6\gamma\,H_{0}^{2}\left(\tfrac{Q}{Q_{0}}\right)^{n}$ model incorporating with two different forms of interacting DE and DM: $3\alpha H\rho_{m}$ and $\frac{\alpha}{3H}\rho_{m}\rho_{DE}$.
+The parameter $\alpha$ in the interacting terms plays a crucial role in determining the viable conditions and estimating the transition from the matter-dominated era (saddle point) to the DE-dominated era (stable node) in viable gravity models at late times.
 As a result, we have discovered fixed points that can be represented as the late-time accelerating universe in $f(Q)$ gravity.
 For the form of ${\cal U}=3\alpha H\rho_{m}$, we have illustrated the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$.
-For $n=3/2$ and $\alpha=1/2$ and $\alpha>0$, we found that the coupling term $\mathcal{U}>0$, signifying energy transfer from dark matter to dark energy implying that the universe was dominated by matter in the early stages and would be dominated by dark energy in later stages.
+For $n=3/2$ and $\alpha=1/2$ and $\alpha>0$, we found that the coupling term $\mathcal{U}>0$, signifying energy transfer from DM to DE implying that the universe was dominated by matter in the early stages and would be dominated by DE in later stages.
 With the current data, fixed point $P_{1}$ is stable and represents the de Sitter acceleration solution, while fixed point $P_{2}$ is a saddle node, and $P_{3}$ is an unstable node that cannot demonstrate universal acceleration.
 Moreover, we have considered another situation of which $n=3/2$ and $\alpha=-2$, $\alpha<0$.
-We discovered for the coupling term $\mathcal{U}<0$ that energy transfers from dark energy to dark matter.
-In this case, the universe was dominated by dark energy in the early stages and will be dominated by dark matter in later stages.
+We discovered for the coupling term $\mathcal{U}<0$ that energy transfers from DE to DM.
+In this case, the universe was dominated by DE in the early stages and will be dominated by DM in later stages.
 With the current data, fixed point $P_{3}$ is stable and exhibits acceleration for the early and present universe but fails to show acceleration for late times, while fixed point $P_{1}$ is a saddle node, and $P_{2}$ is an unstable node.
 For the form of ${\cal U}=\frac{\alpha}{3H}\rho_{m}\rho_{DE}$, the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$ have been examined.
-We have considered $n=3/2$ and $\alpha=4$ ($\alpha>0$) and found that the coupling term $\mathcal{U}>0$, signifying energy transfer from dark matter to dark energy.
-Our results show that the universe was dominated by matter in the early stages and would be dominated by dark energy in later stages.
+We have considered $n=3/2$ and $\alpha=4$ ($\alpha>0$) and found that the coupling term $\mathcal{U}>0$, signifying energy transfer from DM to DE.
+Our results show that the universe was dominated by matter in the early stages and would be dominated by DE in later stages.
 Using the observational data, the fixed points $R_{1}$ and $R_{4}$ were stable and would represent the de Sitter and quintessence acceleration solutions.
-Additionally, for $n=3/2$ and $\alpha=-4$ ($\alpha<0$), we found that the coupling term $\mathcal{U}<0$, meaning energy transfers from dark energy to dark matter implying that the universe was dominated by dark energy in the early stages and would be dominated by dark matter in later stages.
+Additionally, for $n=3/2$ and $\alpha=-4$ ($\alpha<0$), we found that the coupling term $\mathcal{U}<0$, meaning energy transfers from DE to DM implying that the universe was dominated by DE in the early stages and would be dominated by DM in later stages.
 Employing the current data, the fixed points $R_{4}$ and $R_{5}$ have imaginary values, and the remaining point cannot exhibit stable behavior.
-Although many viable $f(Q)$ models explain the dark energy problem in cosmology, our qualitative results from this work can serve as guidelines for more detailed studies.
-They can also be used as complementary constraints on viable $f(Q)$ models, alongside other cosmological constraints on $f(Q)$ theories.
-Our framework, based on the cosmological dynamics of interacting dark energy and dark matter in $f(Q)$ gravity, constitutes the natural template for beyond the standard gravity model, for example, Teleparallel gravity and Gauss-Bonnet gravity, or even more generalizations.
+Although many viable $f(Q)$ models explain the DE problem in cosmology, our qualitative results from this work can serve as guidelines for more detailed studies.
+They can also be complementary constraints on viable $f(Q)$ models alongside other cosmological constraints on $f(Q)$ theories.
+Our framework, based on the cosmological dynamics of interacting DE and DM in $f(Q)$ gravity, constitutes the natural template beyond the standard gravity model, for example, Teleparallel gravity and Gauss-Bonnet gravity, or even more generalizations.
+Advancing our understanding requires developing new theoretical tools to analyze how interactions influence linear and non-linear regimes.
+Precisely determining an interaction kernel is crucial, as it could offer profound insights into fundamental physics, particularly the nature and properties of DM and DE.
 Data Availability Statement
@@
 No.: 201610122060).
-PKS acknowledges Science and Engineering Research Board, Department of Science and Technology, Government of India for financial support to carry out Research project No.: CRG/2022/001847 and IUCAA, Pune, India for providing support through the visiting Associateship program.
+SA acknowledges the Japan Society for the Promotion of Science (JSPS) for providing a postdoctoral fellowship during 2024-2026 (JSPS ID No.: P24318).
+This work of SA is supported by the JSPS KAKENHI Grant (Number: 24KF0229).
+PKS acknowledges Anusandhan National Research Foundation, Department of Science and Technology, Government of India for financial support to carry out Research project No.: CRG/2022/001847 and IUCAA, Pune, India for providing support through the visiting Associateship program.
+We are very much grateful to the honorable referee and to the editor for the illuminating suggestions that have significantly improved our work in terms of research quality, and presentation.
 Appendix A Appendix